### v1

# A Survey on Computational Solutions for Reconstructing Complete Objects by Reassembling Their Fractured Parts

Abstract
Reconstructing a complete object from its parts is a fundamental problem in many scientific domains. The purpose of this article is to provide a systematic survey on this topic. The reassembly problem requires understanding the attributes of individual pieces and establishing matches between different pieces. Many approaches also model priors of the underlying complete object. Existing approaches are tightly connected problems of shape segmentation, shape matching, and learning shape priors. We provide existing algorithms in this context and emphasize their similarities and differences to general-purpose approaches. We also survey the trends from early non-deep learning approaches to more recent deep learning approaches. In addition to algorithms, this survey will also describe existing datasets, open-source software packages, and applications. To the best of our knowledge, this is the first comprehensive survey on this topic in computer graphics.
Authors contributed equally.  Corresponding authors.
Keywords Computing methodologies $\cdot$ Computer graphics $\cdot$ Shape modeling $\cdot$ Shape analysis
1 Introduction
Reconstructing a complete object from its parts is a fundamental problem in many scientific domains. In archaeology, we typically collect fragments of one or multiple objects at an archaeological site, and the goal is to restore the original objects. When people have bone fractures, we need to study how to assemble bone fragments to plan bone reduction surgery. This problem also arises in the 3D object/scene modeling, where we want to construct a complete object/scene from parts/objects. Other relevant problems are protein docking and human object interactions, in which the difference is that the transformation of each part becomes non-rigid. The purpose of this article is to provide a systematic survey on this topic, including algorithms, datasets, open source software packages, and applications. To the best of our knowledge, this is the first systematic survey on this topic.
The reassembly problem is related to many shape analysis tasks. An important task is to segment the regions of the fractured surfaces and the original surfaces, which is a special shape segmentation problem. Another important task is to find correspondences between fractured regions of different pieces to find relative poses between them. This is related to scan matching, with the difference in that the correspondences are between complementary points. In addition, many approaches utilize priors of the underlying complete object, including symmetries and shape priors. This calls for approaches to analyze and encode such priors and use them to solve the reassembly problem.
Similar to other problems in computer graphics, we have also witnessed the transitions from early non-learning based approaches to recent deep learning approaches. Early approaches have explored all aspects of the reassembly problem, including segmentation of fractured regions and original regions, matching fractured regions, and encoding priors (symmetries and shape spaces) of the underlying complete shape. However, they are restricted to hand-crafted features, so the entire systems do not generalize robustly. In addition, it is difficult to encode shape priors among objects with large geometrical and topological variations. These issues can be addressed using deep learning approaches. Besides the fact that the shape segmentation and shape matching problems become more robust, we also have approaches that perform shape segmentation and shape matching together, which benefit from end-to-end training in deep learning. In addition, with advances in learning 3D shape generative models, it is possible to learn shape priors from diverse shape collections to reassembly.
Figure 1: Overview of this survey. (a) Single-piece processing and analysis. (b) Multi-piece analysis. (c) Template shape space and template-based reassembly. (d) Applications.
Despite tight connections to relevant problems in shape analysis, applying existing approaches to the 3D reassembly problem is not straightforward. This is the case for both non-deep learning and deep learning based approaches. However, in the literature, people do not establish the reassembly as a unique sub-area in geometry processing and shape analysis. Our goal is to summarize state-of-the-art results in this domain systematically, seeking to establish the skeleton of this important sub-area.
1.1 Related Surveys
The most relevant survey is geometric analysis in cultural heritage [1], in which reassembling fractured objects is a central topic. The difference between our survey and [1] is two-fold. First, in contrast to their survey which focuses mostly on non-deep learning techniques, our survey covers recent advances using deep learning. Second, in contrast to the specific application of assembling fractured objects, this survey covers many applications, including bone fracture reduction, shape modeling by assembling parts, and applications in biology, geoscience, and paleontology.
Another relevant survey is [2], which studies how to decompose a shape into rigid parts. Its focus is different in the sense that it focuses on optimizing decompositions with respect to different geometric and physical criteria for manufacturing and fabrication. In contrast, we focus on solving the inverse problem of reconstructing complete objects from parts.
Our survey also discusses geometry processing techniques, including 3D reconstruction [3, 4], mesh segmentation [5, 6], primitive segmentation [7], shape descriptors [8] and correspondences [9], symmetry detection [10], and data-driven shape analysis and processing [11]. While the pipeline of reassembling parts uses these techniques, this article focuses on approaches tailored for this particular class of applications, which usually have modified inputs, outputs, and problem setups.
Many reassembly techniques utilize shape priors of the underlying complete priors. This is an area in which we have seen great progress in adopting machine learning models to learn shape priors from data under suitable 3D representations. Recent surveys [12, 13, 14] provide excellent coverage of state-of-the-art results. In contrast, this survey focuses on how to use such parametric shape priors to solve the reassembly problem.
There are also other relevant review articles. [15] surveys related approaches for each step in the pipeline that starts from reassembly to object completion, which also addresses missing regions. [16] present a survey article on the specific problem of reassembly of pottery fragments. [17] surveys related work on reassembly of defective cultural heritage objects. Our survey presents a comprehensive review of general fractured object reassembly and connections to relevant research results in geometry processing.
1.2 Paper Organization
We organize the reminder of this paper as follows. In Section 2, we provide a more detailed overview of this paper. In Section 3, we describe existing approaches for analyzing individual pieces. In Section 4, we discuss existing approaches for analyzing multiple pieces, focusing on problems of matching pairs of pieces and multiple pieces jointly. In Section 5, we introduce existing methods for encoding prior knowledge of the underlying complete object. In Section 6, we review approaches that utilize priors about the complete object to solve the reassembly problem. In Section 7, we discuss various applications of the reassembly problem. Section 8 present available datasets and software packages. In Section 9, we present current trends and future directions on this topic. Finally, we conclude this survey in Section 10.
2 Overview
This section presents an overview of this survey (See Figure 1).
Single piece processing and analysis (Sec. 3). Single piece processing and analysis concerns individual pieces of fractured objects. Topics include how to digitize physical fragments into digital 3D models and how to segment each fragment into regions of original surfaces and fractured surfaces. Although these topics are central topics of the field of geometry processing and shape analysis, we focus on techniques that are tailored for the particular characteristics of fragments of physical objects. We will cover both non-deep learning approaches and more recent deep learning approaches.
Multi-piece analysis (Sec. 4). Multi-piece analysis studies the critical problem of establishing correspondences between fragments to match them. We begin with pairwise matching techniques using geometric and textual features extracted from the fracture surfaces and original surfaces. We will survey global techniques that do not require prior knowledge of the relative pose between a pair of fragments and local techniques that start from reasonable initial relative poses. We then proceed to discuss methods that involve more than two pieces, in which there is a cycle-consistency constraint among the relative poses along a cycle of pieces. In addition to fragment matching, we will also discuss approaches that solve the correlated problems of fragment segmentation and fragment matching together. Similarly to single-piece analysis, we will cover both non-deep learning approaches and deep learning approaches.
Template shape space (Sec. 5). Template analysis studies how to encode prior knowledge of the underlying complete object. We focus on techniques that learn the shape spaces of complete objects from the data. Early approaches of this type focus on learning linear shape spaces. Recently, we have seen great progress in adopting generative modeling techniques to learn more complex shape spaces. Challenges on how to apply the learned shape priors for fractured object reassembly are discussed.
Template-based reassembly (Sec. 6). We then discuss approaches that used priors of the underlying complete objects to solve the reassembly problem. These include approaches that enforce priors of symmetries and primitive structures and those that optimize a complete object in the learned shape space. In the latter case, a central problem is to establish correspondences between points on the original surfaces of the fragments and points on the template surface.
Applications (Sec. 7). After discussing technical approaches towards solving the reassembly problem, we will survey various applications of the reassembly problem. We will discuss five areas of fractured object reassembly application, including bone reduction, restoring fragments of thin-shells, walls, and building, and reassembly problems in science. Finally, we discuss the relevant problem of assembling parts to form complete objects and assembling objects to form 3D scenes.
Datasets and software packages (Sec. 8). We discuss publicly available datasets and open source software packages. Datasets include early small-scale datasets and recent large-scale real- and synthetic datasets.
Future directions (Sec. 9). The last paper of this survey summarizes existing approaches and discusses future directions. We will emphasize how to advance existing reassembly problems by developing better learning approaches, focusing on challenges in representations, data, and utilizing foundation models. Moreover, we will also discuss potential new applications of reassembly in natural and life sciences.
3 Single Piece Processing and Analysis
This section presents techniques for single-piece analysis. We begin by examining relevant work on digitizing fragment shapes in Section 3.1. We then discuss prior work on segmenting a piece into fracture surfaces and original surfaces in Section 3.2.
Table 1: Category of scanning technology.
Categories Accuracy Efficiency Cost(\$) Key features
Laser $0.01\sim 3$ mm several minites $500$ $5$k Fast capture of large pieces
Photogarmmetry $<=0.15$ mm minutes to hours $200$ $5$k Capture of color and texture information
Structured light $0.1$ mm seconds to minutes $50$ $5$k High efficiency and precision
CT $0.001\sim 0.5$ mm minutes to hours $1$k - $10$k Internal structure detail capture
3.1 Reconstruction
Physical fragments must be captured and reconstructed into digital models before reassembling. This section discusses how fragmented pieces can be captured using various sensors, each requiring specific reconstruction techniques.
Laser scanning. As shown in Figure 2 (Top-Left), laser scanners capture the 3D geometry of fragments by emitting laser beams and measuring the time or phase difference of the reflected light. They allow for high-resolution geometry capture, achieving a physical distance of approximately $0.01$ mm between each captured point. The acquisition efficiency for capturing a piece, depending on the resolution selected, laser emission frequency, and the number of emitted laser lines, typically ranges from minutes to hours. For example, scanning a $10$ cm object using the low-frequency NextEngine scanner ($50K$ points per second) would take about an hour [18]. However, with high-frequency HandySCAN ($1.8M$ points per second), the process would only take a few minutes [19, 20]. Laser scanning is usually desirable for capturing high-quality geometry, particularly in outdoor environments. However, since it does not capture color information, other capturing modalities should be combined if chromatic data is needed. A significant drawback of these scanners is their high cost, though rental options are available. There are several representative datasets captured by laser scanning. Stanford University used laser scanners to collect the “Forma Urbis Romae” dataset as part of the Digital Michelangelo Project, capturing high-resolution 3D shapes of marble fragments for digital preservation and research [21]. Another significant dataset is generated through the “Reconstructing Thera Frescoes” project, where low-power laser scanning is employed to ensure accurate data capture while also protecting fragile frescoes from potential damage during the scanning process.[22]
Photogrammetry capturing. Color images captured by regular cameras can also be used to capture fragment geometry and texture; and recent studies have found that geometry reconstruction quality can be comparable to laser scanning, with a resolution of $0.15$  mm and an accuracy of $0.13$  mm [23, 24]. State-of-the-art multi-view stereo (MVS) algorithms reach pixel level accuracy, namely, $90\%$ reconstructed points have accuracy within about one pixel [25]. Therefore, the geometry quality can be further improved by capturing detailed images from a short range. As shown in Figure 2 (Top-Right), photogrammetry is well suited for field data collection because it does not require specialized skills to operate a scanner, is affordable and easy to set up. Designed for field collection without electricity, [26] proposed a rig to do photogrammetrical documentation of artifacts. Data collection for this protocol averaged approximately $20$ minutes, with a series of overlapping photographs (approximately $100$ per artifact) to be taken of the artifact from all angles [27]. To obtain even higher throughput, the batch method has been proposed by [24], achieving a throughput of about $730$ fragments per day. However, photogrammetry is sensitive to light and is not suitable for transparent material. The lighting needs to be closely controlled using two overhead fluorescent lamps or LEDs, which minimize the amount of shadows and regularized the amount of light across the surface[18]. Note that a list of “non-friendly” materials for photogrammetric purposes, such as water, glass, varnish, steel, and many more, can be found in [28]. In addition, selecting an appropriate focal length is essential to achieve optimal results in the digitalization process [29]. Wang et al. proposed a comprehensive dataset consisting of batches of 3D sherd scans, typically pottery fragments, accompanied by auxiliary information that aids in the alignment and reconstruction of the original artifacts through model registration techniques [24].
Figure 2: Four types of sensors for digitizing fragment shapes. (Top-left) Laser scanning [21]. (Top-right) Photogrammetry capturing [24]. (Bottom-left) Structured light capturing [30]. (Bottom-right) CT Scanning [31].
Structured light capturing. As shown in Figure 2 (Bottom-Left), The main advantages of this technology are speed, making it suitable for applications where efficiency is crucial [32]. This method operates by projecting structured light onto an object and then analyzing the reflected light to determine the shape and dimensions of the object. The resolution of structured light scanners, which is around $0.1\sim 0.2$ mm, is slightly inferior to that of laser scanners. However, it can generally obtain better geometric detail and resolution in small details than digital photogrammetry. The operation time can range from 1 to 10 minutes, depending on the optical properties of the scanned object. In [33], 60 pottery fragments can be scanned per hour at the desired resolution. For some reflective pieces, such as flint, an experienced operator produced complete digital 3D images at the rate of $5\sim 6$ artifacts per hour [34]. Compared with laser, structured light can capture color information, but may struggle with shiny or dark surfaces. Therefore, they are typically well-suited for scanning small parts that require color and texture details, and for use in indoor environments. A representative project in this category is [35].
CT scanning. CT scanners are extremely useful in piece reconstruction because they allow 3D scanning not only of the outer surfaces of an object but also hidden internal structures [36, 37]. Unlike outer surface geometry, internal structures of the pieces are difficult to capture using the aforementioned scanning technology that relies on light reflection. The resolution is typically approximately $1\sim 500$ microns, a level of precision capable of capturing the fine topology [38]. To date, numerous studies have used CT to image rocks [39], stones [40], fossils [41], bone [42] and teeth [43]. Even though a high-resolution CT-scan can take several hours, the duration needed to scan a single part is manageable. [40] were able to scan up to 220 lithics at once in a relatively short period of time (ca. 2 hours) at a resolution of $140$ microns. Their tests suggested that this resolution is optimal for scanning many artifacts simultaneously while maintaining accuracy in the 3D shapes. The skull dataset [44] comprises high-resolution CT scans of both damaged and normal skulls, aiding in the study and reassembly of cranial anatomy. Another dataset [45] consists of high-resolution CT scans generated through helical scanning, providing detailed 3D reconstructions of anatomical structures stored in DICOM format.
Registration and 3D reconstruction. In order to obtain the complete geometry of an object, it is necessary to scan from multiple angles and then use point cloud registration to align these scans in a common coordinate system. When reference markers are possible, such as small disc-shaped stickers [19] or unevenly spaced grid lines drawn on a turntable [18], they facilitate point cloud registration by serving as common reference points. When using such markers is infeasible, markerless registration such as Iterative Closest Point (ICP) algorithm and its variants are utilized to align the point clouds [33]. The quality of the initial setup in the Iterative Closest Point (ICP) algorithm is crucial, as it significantly influences the convergence rate and precision of the algorithm [46]. Therefore, careful manual annotations are essential for the optimum outcomes. In order to improve registration efficiency, [24] devised an automated and robust registration technique that leverages the bilateral boundary ICP, offering reliable initial relative pose estimates to initiate the ICP. After forming the dense point cloud, a mesh modeling algorithm, such as marching cubes [47] or Poisson reconstruction [48] is used to generate the surface, In contrast to the methods that record the scans before reconstructing the surface, [49] integrates both processes simultaneously, resulting in a noise-free and watertight solid model. Using joint registration, the method significantly improves reconstruction accuracy, ensuring better consistency when combining different views. The experimental results indicate that it can perfectly reconstruct the surfaces of archaeological fragments. This approach was used to reconstruct the fragments of the Octogan project [50].
Batch processing in scanning improves efficiency and reduces the need for manual intervention. This approach is widely adopted in large-scale applications. [35] proposed a system to automatically scan multiple 3D objects using view and path planning algorithms, minimizing manual intervention. It enhances scanning efficiency and quality with structured light scanning and calibrated positioning, making it ideal for large-scale applications like archaeological documentation. [22] introduced a batch processing system that achieves efficient scanning through parallel operation of multiple devices, automated processing, and a fast matching algorithm. A single operator can manage up to four 3D scanners, while another uses a flatbed scanner to capture high-resolution images and normal maps. The system automatically aligns and processes data in the background, minimizing manual intervention. The fast geometry-based matching further optimizes efficiency, allowing the system to process approximately 10 fragments per hour. [24] presented another batch system to scan two sides of fragments and developed tailor registration and reconstruction techniques to process the paired scans.
3.2 Segmentation
Surface segmentation is a customary preprocessing step in the analysis of individual fragments for reassembly tasks, aimed at breaking down complex surface geometries into simpler, meaningful regions. Typically, this involves distinguishing between the original and fracture surfaces, facilitating easier analysis and subsequent matching with other fragments to achieve successful reassembly.
Geometry-based method. Geometry-based surface segmentation methods rely on the intrinsic geometric properties of the fragment’s surface, such as curvature, edges, and other spatial features. These methods are among the most established techniques in the field due to their reliance on fundamental geometric principles (see Figure 3 (Top)). Several geometry-based methods have been developed for surface segmentation. For instance, [51] employ a simple region-growing algorithm primarily based on normal vectors, along with a facet segmentation threshold, to initially segment the original object into distinct surface regions. Following this segmentation, bumpiness is used to detect fractured surfaces. [52] first perform multi-scale edge extraction on point-sampled surfaces, which is constrained to return closed cycles to obtain initial face cycles. These face cycles are then partitioned into original and fracture surfaces using an iterative normalized cut method on a weighted graph, constructed considering surface sharpness and roughness. [53] use a hierarchical agglomerative clustering algorithm that starts with each element as its own cluster, merging the closest clusters based on a user-defined metric until no further merges are possible. A custom 2-level caching scheme is used to reduce redundant distance calculations and sorting, enhancing efficiency. [54] first apply the Laplace operator to smooth the surface, and then use a clustering algorithm based on a vertex normal vector to acquire rough surface segmentation, remove noise surfaces, and then refine the segmentation by merging adjacent faces using surface roughness and face normal vectors. In addition to these, various other clustering and graph segmentation algorithms, such as K-means clustering [55], hierarchical clustering [56], mean-shift clustering [57], spectral clustering [58] and random walks [59], can also be adapted for fractured surface segmentation. These techniques are well-established, relying on explicit rules and geometric properties. They tend to be more interpretable and less data-intensive but are limited by surface complexity and noise.
Figure 3: (Top) Non-deep learning methods for fracture surface segmentation [51]. (Bottom) Deep learning methods for fracture surface segmentation [60].
Figure 4: A representative multi-piece shape analysis framework presented in [52]. Pair-wise matching use use patch features extracted from fracture surface and segment features extracted from fracture surface boundaries. A voting-based approach is used to find consistent correspondences between geometric features. Matching among pieces employ a greedy approach.
Learning-based method. Learning-based surface segmentation techniques have gained traction in recent years due to their ability to automatically learn patterns and features from data, without the need for explicit geometric rules. These methods, particularly those involving deep learning, are transforming how surface segmentation is performed in fragment reassembly tasks. Although some learning-based fractured object reassembly methods take the entire fragments as input for analysis, bypassing prior surface segmentation, there are still methods that decouple the reassembly pipeline, retaining segmentation to achieve better matching and reduce computational costs. Learning-based segmentation methods are predominantly based on graph learning techniques. The initial segmentation process can be conceptualized as a binary classification task. For example, [60] utilize a convolutional neural network (CNN) to learn the labels of ridge and non-ridge cells (see Figure 3(Bottom)), facilitating the extraction of the fracture margins from the fragments. Subsequently, fractured surface can be well segmented by trimming the fragment models along these margins. [61] employ Multi-Layer Perceptron (MLP) layers combined with a sigmoid activation function to predict the confidence in identifying a fracture point. If we generalize segmenting fractured surfaces to 3D semantic segmentation, we can also leverage advanced general-purpose graph learning algorithms like PointNet [62], PointNet++ [63] and DGCNN [64] to perform segmentation. These powerful frameworks offer robust solutions for intricate 3D data analysis, and they will also be mentioned in the next section. These approaches, particularly deep learning, offer more flexibility and can handle complex surfaces and noisy data better. However, they require large datasets and significant computational power.
4 Multi-Piece Analysis
This section reviews the literature on multi-piece analysis, which considers two or more fragments. In Section 4.1, we review pairwise matching techniques. In Section 4.2, we discuss multi-piece matching techniques, including early optimization-based techniques and recent techniques based on deep pose regression. Finally, we discuss the correlated problem of joint segmentation and matching in Section 4.3.
4.1 Pairwise Matching
Rigid scan maching. A relevant problem of pairwise piece matching is rigid scan matching, which finds the relative rigid pose between two depth scans. Rigid scan matching is an important step in 3D geometry reconstruction from depth scans. A standard pipeline of rigid scan matching consists of feature extraction, feature matching, and rigid pose fitting. Early approaches use hand-crafted geometric descriptors, e.g., SpinImage [65] and Integral invariant [66] features. These features usually only encode local information; therefore, the resulting feature matches have a lot of false positives. The rigid pose fitting procedure addresses this issue by enforcing the rigidity constraint, which finds a subset of feature matches that share a rigid transformation. Popular rigid pose fitting approaches include RANSAC [67], general Hough transform [68, 69, 70], and their variants [71, 72]. Another category of methods formulates rigid scan matching as spectral matching, which computes a consistency matrix among candidate feature matches. The elements of this consistency matrix encode whether distances and angles are preserved between the corresponding pairs of correspondences. Rigid pose fitting is formulated as computing a sub-matrix in which all elements are consistent, and this is done using spectral techniques.
More recent approaches develop deep learning techniques for rigid scan matching. These approaches employ deep neural networks to extract features and develop differentiable layers for fitting rigid poses. Examples include differentiable RANSAC and attention-based rigid regression. Representative methods in this category include DCP [73], PNetLK [74], FCGF [75], Predator [76], RoReg [77], and GeoTF [78]. These approaches show significantly improved performance on benchmark datasets where the domain gap between the testing set and the training set is small. However, it is still an open question to develop generalizable deep learning approaches for rigid scan matching.
Non deep fragment matching. Similar approaches have been developed for rigid fragment matching, which also combines feature extraction, feature matching, and rigid pose fitting. However, rigid fragment matching and rigid scan matching have several notable differences. These differences come from several aspects, including where to extract features and what constraints are induced from feature matching. Another important difference is the robustness of the extracted features, the feature descriptors, and the matching procedure, as the surfaces of the fragments may be severely damaged or there are tiny fragments between two fragments, causing geometric differences. Figure 4 shows a representative framework for fragment matching.
Figure 5: Matching features on original surfaces. (Left) Stanford project [79]. (Right) The idea of matching completed objects with fragment extrapolation [80].
Most approaches use features extracted from the fracture surfaces. This requires fracture surface segmentation as discussed in Section 3.2. Robustness has been a major concern when extracting features from fracture surfaces. Therefore, many features introduced in the literature are patch features that are introduced to improve robustness. [52] use patches defined by sub-level sets of the integral invariant descriptors and summarize geometric descriptors derived from performing principal component analysis of patch shapes. [81] introduced an approach to smooth the fragments and segment the concave and convex regions of the fracture surfaces for robust matching. [82] studied multi-scale covariance matrix based point descriptors. [83] introduced an algorithm based on cluster-tree tailored for matching fragments with salient fractured surfaces. A key issue in matching fracture surfaces is developing descriptors that can be used to identify geometric patches that can be tucked together [84]. A recent paper [61] also addresses this issue by introducing a primal-dual descriptor in deep learning approach.
Many papers focus on extract from boundary curves of fracture surfaces [84, 81]. Such features are more robust in the setting where the areas of the fracture surfaces are small, making it difficult to compute and extract geometric descriptors. [85] analyze and simulate fracture patterns of Theran wall paintings. It builds statistical models of boundary curves of wall painting fragments, which can be applied to match boundary curves. [84] detect boundary curves based on curvature computation and apply FFT to compute the descriptors of points on the boundary curves for fragment matching. [86] introduced a multi-channel boundary contour representation for matching sub-contours via 2D partial image registration.[87] perform curvature computation to detect parabolic contours, whose descriptors are determined by moments. [88] studied how to detect boundary curves of thin fragments. Fragment matching is done using voting. [89] introduced a fast algorithm for 2D fragment assembly based on a partial EMD distance. [90] introduced a descriptor based on the convex/concave information of a point on the boudnary curves of the fractured surface. The similarity between two fractured surfaces is calculated based on the spin images of the point of the feature curve and the distance and normal deviation between the two feature curves to find the matching fractured surface. [91] utilized the detected breaking curves as cues for fracture surface segmentation and registration. [92] introduced a new matching algorithm for stone tool reassembly based on detecting and matching contour points on flake surfaces. Matching is done using a five-point method, which extends Super4PCS. For efficiency concerns, it focuses on matching flake surfaces to a core stone, which is applicable in this specific setting. In [93], a set of matching units is detected and described by the 2D Link-Chain Descriptors (LCD) and the 3D Spatial-Distribution Descriptors (SDD). Second, the pairwise reassembly probability is calculated using LCD and SDD descriptors; then, collision detection is performed to eliminate the incorrect overlapping pairs.
Several approaches rely on features on the original surfaces for matching (see Figure 5). Early work by [94] discussed how to use original surface features for geometric matching. Examples include the continuity of texture patterns or matches between curve features extracted from original surfaces. [86] introduced a similar approach to enforce the original surface consistency. The former approach can be found in most reassembly pipelines of wall painting fragments, e.g., [79]. [80] studied how to match fragments that are under erosion and proposed an approach that completes the original surfaces of the fragments and matches the completed surfaces. In this approach, the original non-overlapping surfaces become overlapping, allowing us to detect and match extracted feature points.
Figure 6: Robust scoring functions to evaluate the quality of a match proposed in [95]. (Top-left) Criteria 1, gap volume. (Top-right) Criteria 2, contact area. (Bottom) Criteria 3 and 4, bounds for the area and the mean curvature, and the length of contact curves.
In many cases, the approach of extracting and matching features is not robust when fragments are undergoing serious erosion, which is the case for most real datasets. In [96], the authors find that detecting and matching feature points from fractured surfaces is much more challenging than matching scans. They proposed confidence scores on potential matches. Several approaches define scoring functions over the space of transformations and develop global search strategies to global optimal solutions. [97] introduced an approach that focuses on wall-painting fragments and applies a discrete search of relative poses to identify global solutions to reassembly. [95] introduced four criteria for fragment matching (see Figure 6). The first criterion exploits the volume of the gap between two properly placed fragments. The second considers the overlap of the fragments in each possible matching position. Criteria 3 and 4 employ principles from the calculus of variations to obtain bounds for the area and the mean curvature of the contact surfaces and the length of contact curves, which must hold if the two fragments match.
In addition to geometric matching, another difference between scan matching and fragment matching is that the matches should be free of penetrations. This constraint is easy to integrate for RANSAC-based approaches [67, 98]. An open question is how to enforce this under global optimization approaches such as spectral matching. Regarding local refinement methods, several papers [52, 99, 100, 101] studied how to formulate constrained optimization techniques to enforce non-penetration constraints (See Figure 7 (Left)). [50] introduced a gradient field to optimize the packing of a collection of objects, enforcing the non-penetration constraint (See Figure 7 (Right)).
Finally, several systems employ users in the loop and develop semi-automatic approaches for fractured object reassembly. In [102], user input is used to prune incorrect matches.
Deep fragment matching. In the deep learning era, the research community has studied developing deep neural networks to solve the multi-piece reassembly problem. Early work includes Deepzzle [103], which studied image reassembly, a special case of fractured object reassembly. Deepzzle studied how to handle noisy inputs, e.g., erosion, missing pieces, and fragments of other objects. Thanks to the emergence of 3D benchmark datasets for fractured object reassembly, there are increasing interests in developing deep learning approaches to solve the 3D fractured object reassembly. [104] proposed MatchMakerNet (see Figure 9(Left)), a network architecture designed to automate the pairing of object fragments for reassembly. The approach takes two point clouds as input and its network leverages graph convolution alongside a simplified version of DGCNN for prediction.
However, most existing approaches have focused on feature learning and incorporating rigidity constraints and problem-specific knowledge deserves future research.
Figure 7: Enforcing non penetration via constrained optimization [52, 99, 100, 101].
4.2 Multi-Piece Matching
Optimization-based based approaches. The difference between multiple-piece matching and pairwise matching is that there is a cycle-consistency constraint [105, 106, 107] among cycles of fragments. Specifically, when composing relative poses along a cycle of fragments, we should arrive at the identity map. If not, then at least one of the relative poses is incorrect. This cycle-consistency constraint can be used to prune incorrect pair-wise matches computed between pairs of objects in isolation.
In the literature on scan matching, early approaches enforce combinatorial optimization to search for consistent matches. A popular way is to optimize a spanning tree among the input graph of pairwise matches [108, 109, 52]. The advantage is that it is easy to incorporate the non-penetration constraint among fragments. For example, in the case of scan matching Huber and Herbert [109] showed how to incorporate the visibility constraint in multi-scan matching. On the other hand, combinatorial optimization approaches are typically greedy and can easily return sub-optimal solutions.
Another category of approaches formulate multi-piece matching as Markov-random field (MRF) inference. This is formulated as sampling a discrete set of candidate poses of each object and model pairwise matches as pairwise potentials to optimize the best candidate of each object. This approach was first introduced to solve the Jigsaw puzzle problem in images [110, 111], where the candidate positions of each image patch are given by the grid of image patches of a given image. A relevant approach is [112], which presents a linear programming (LP) relaxation to the MRF formulation on the image-based square patch puzzle problem. This MRF formulation was later applied for in MRF-SFM [113, 114] to solve the structure-from-motion problem. The nice properties of this MRF formulation is that it possesses better solvers that offer improved solutions compared to greedy approaches. Moreover, it is easy to incorporate the non-penetration constraint, i.e., if the a pair of fragments penetration under two candidate poses, then the corresponding pairwise potential is infinite. The limitation of this MRF formulation is that it requires many pose samples to obtain accurate solutions.
Modern multi-piece matching typically formulates low-rank matrix recovery by establishing the equivalence between the cycle-consistency constraint and the fact that the matrix that encodes pairwise maps in block is low-rank, c.f. [115]. The optimization approaches fall into categories of semidefinite programming relaxation [116, 117, 115, 118, 119, 120, 121], spectral techniques [122, 123, 107, 124, 125, 126, 127, 128, 129, 130, 131, 132], and non-convex optimization [133, 134, 135, 136, 137, 138, 139]. Their advantages are superior empirical performance and theoretical guarantees. However, it is not easy to enforce the penetration-free constraints under them.
Figure 8: Multi-fragment matching [140], which enforces the consistency of matching along loops.
Figure 9: Recent advances on deep fragment matching. (Left) Techniques that extend rigid matching [104]. (Middle) Transformer-based techniques [141]. (Right) Equivariant techniques [142].
Applications of these multi-object matching approaches in the context multi-fragment reassembly have been limited. Most approaches follow greedy approaches [52, 143, 141]. There are a few exceptions. [144] introduced an unsupervised genetic algorithm to reassemble wall painting fragments. The approach evolves a pool of partial reconstructions that grow through recombination and selection over the course of generations. The authors introduced a novel algorithm for combining partial reconstructions that is robust to noise and outliers, and a new selection procedure that balances fitness and diversity in the population. [86] detect loop in the pose graph and performs loop closure to rectify pose errors. [140] enforce global Consistency in the Automatic assembly of fragmented artifacts. This approach addresses error accumulation problem the global reassembly of many fragments, among which there are many potential matches with a pair of fragments. [93] introduced a global optimal reassembly solution is obtained by iterative graph optimization with the constrains of the loop-closures and overlap restrictions. One critical difference between multi-scan matching and multi-fragment matching is that the underlying graph of the fragments is much more sparse than the graph in multi-scan matching, as scans have significantly overlap, while fragments do not overlap.
Deep pose regression based approaches. The emergence of large-scale datasets such as Breaking Bad [145] has simulated a wave in developing deep learning approaches that perform pose regression. Many approaches adopt recent advances in 3D feature learning using Transformers. [146] developed Transfomers to find pairs of matching fragments and their relations. Specifically, they first employ a graph-based network to extract the local contour and texture features of the fragments. Then, for the pair-searching task, they adopt a linear transformer-based module to integrate these local features and use contrastive loss to encode the global features of each fragment. For the pair-matching task, they design a weighted fusion module to dynamically fuse extracted local contour and texture features, and formulate a similarity matrix for each pair of fragments to calculate the matching score and infer the adjacent segment of contours. [147] introduced Proxy Match Transform (PMT), an approximate high-order feature transform layer that enables reliable correspondences between the mating surfaces of parts while being efficient. Using PMT. they introduced a new framework, dubbed Proxy Match TransformeR (PMTR), for the geometric assembly task. [148] introduced a Transformer-based encoder to extract both global and local geometric features, which are then passed to two decoders to estimate the relative pose between two fragments. [141] introduced a recurrent approach that aligns and merges fragments into larger groups. The key contribution is a diffusion model simultaneously denoises the 6-DoF alignment parameters of the fragments, and a transformer model verifies and merges pairwise alignments into larger ones, whose process is repeated iteratively. [149] introduced an approach that identifies critical stones among the input fragments for efficient computation of the reassembly procedure. This is implemented using the attention mechanism in an iterative manner. [150] use a hybrid attention module to model and reason complex structural relationships between fragment patches. The model has intra- and inter-attention layers, enabling capturing of crucial contextual information within fragments and relative structural knowledge across fragments. The model is enhanced with an adjacency aware hierarchical pose estimator.
People have developed equivalent networks [151] for point cloud learning to solve fractured object reassembly. [142] developed equivariant and invariant encoders of each part to compute equivariant and invariant correlation modules to aggregate extracted signals across all parts, which are used to regress the pose of each input part. [152] introduced an approach, called SE(3)-bi-equivariant transformer (BITR), based on the SE(3)-bi-equivariance prior to the task. Specifically, BITR first extracts features of the inputs using a novel SE(3)×SE(3)-transformer, and then projects the learned feature to group SE(3) as output. BITR can not only handle non-overlapped PCs, but also guarantee robustness against initial positions. Extension of SE(3) equivariance. [153] introduced DiffAssemble, which follows Diffusion Probabilistic models formulations. DiffAssemble models a Markov chain in which it injects noise into the position and orientation of the pieces. The correct and random poses at the initial time step and the final time step, respectively. The denoising network is an attention-based GNN that takes as input a graph where each node contains an equivariant feature that describes a particular piece and its position and orientation. Similarly, [154] and [141] treated the reassembly task as a denoising task and introduced FragmentDiff and PuzzleFusion++, a method that applies diffusion denoising through transformer to predict the pose parameters of each fragment based on their global feature correlations and learned pose priors.
4.3 Joint Segmentation and Matching
Figure 10: Joint matching and segmentation [61].
Figure 11: (a) Early work on linear shape spaces [155]. (b) Deep deformable shape generators [156]. (c) Implicit shape generators [157, 158, 159]. (d) Point cloud shape generators [160].
As discussed above, an important problem in fragment analysis is segmenting the fracture surface. Typically, this is done for each fragment in isolation. However, the segmented surfaces should be compatible with those of adjacent fragments. This means that the segmentation problem and the fragment matching problem are correlated and should be solved together. This joint segmentation and matching problem first appeared in the context of joint shape analysis, in which segments of an object provide a compact representation that relates shapes that are structurally relevant but differ in geometric details. [161, 162] introduced joint optimization approaches to optimize the segmentations of individual parts and the correspondences between the segments. [163, 164, 118, 165] introduced the functional map framework which allows us to optimize consistent segmentations by optimizing a function sub-space of each shape.
In the context of fractured object reassembly, the only approach that explicitly models this objective is JIGSAW [61] (see Figure 10), It extracts point-wise features using transformers, which are used to predict point labels of original or fracture surfaces and perform point-wise matching. Joint segmentation and matching are achieved using shared deep point-wise feature descriptors.
5 Template Shape Space
In many cases, we do not have prior knowledge of the exact shape of the underlying complete object. However, we may have prior knowledge about its category, which provides shape priors. Therefore, an important problem is learning a parametric shape space from a collection of example shapes. This problem has been studied extensively in the literature. This section presents relevant approaches that can be applied for fractured object reassembly. We differentiate approaches that focus on organic shapes with the same topology and man-made shapes with varying topology.
Early approaches (see Figure 11(a)) include Morphable faces [166] and SCAPE [68], which focus on 3D shape spaces with the same topology. Recent examples along this line include SMPL [155] for humans, SMAL [167] for animals, MANO [168] of hands, and 3DMM [169] for faces, which are learned from larger datasets and therefore exhibit significantly improved performance in solving inverse problems. These models decouple shape variations and pose variations. Shape variations are derived from PCA analysis. Pose variations, e.g., in SMAL and SMPL are encoded using bones. A limitation of these approaches is that require consist meshes to learn the shape prior.
Recent approaches have focused on using deep neural networks to learn generative models, which are superior to linear models. Most approaches [170, 171, 172, 173] use graph neural networks to decode a latent code into a mesh. Many approaches focus on network design. ARAPReg [174] and GeoLatent [156] introduced regularization losses that model shape deformations between adjacent synthetic shapes to improve the generalization of the shape generator. Despite significantly improved performance from linear models, these approaches still require consistent inter-shape correspondences as input for learning.
DeepSDF [157] and follow-up work show that shape generative models can be learned from raw data points with correspondences. SALD [175] shows that MLP-based implicit surfaces can also encode deformable shapes, but generalization of shapes that undergo large pose changes is very limited. GenCorres [139] addresses the generalization issue under large pose changes by developing an approach to compute dense correspondences between adjacent implicit surfaces, allowing us to apply ARAPReg to define regularization losses. It leads to significantly improved synthetic shapes under novel poses.
Significant progress has been made on learning generative models to synchronize man-made shapes. Most approaches that produce high-quality synthetic shapes are under broadly defined implicit shape representations (signed distance functions, volumetric grids, and neural fields), These include early efforts [176, 158, 177, 178, 179, 180] under the generative adversial network (GAN) [181] and variational auto-encoder (VAE) [182] paradigms and recent results [183, 184, 185, 159, 186, 187] under the latent diffusion paradigm [188]. However, one limitation of using the implicit representation is that it is very difficult to compute the correspondences between the learned template and the input fragments.
Figure 12: Using prior knowledge from a complete object model to aid reassembly. (a) Local-global alignment with a selected explicit template [189], such as a skull. (b) Implicit learned prior knowledge from a large-scale 3D object dataset [190]. (c) A global evaluation using neural network as complete object prior [191, 192].
Table 2: Category of the different complete object priors with their descriptions and limitations.
Categories Descriptions Limitations
Local-global Alignment Shape prior provides reasonable initial reassembly Applicable to limited types of reassembly objects
Global Evaluation Regularize the quality of the reassembled result Hard to find a proper function for each reassembly
One way to address this issue is to learn generative under explicit representations such as point clouds and meshes. Many approaches have shown promising results for point cloud generators for man-made shapes [193, 194, 195, 160, 196]. Although it is possible to fit a point cloud generator to shape collections with large geometric and topological variations, there are two limitations that prevent them from serving template models for fractured object reassembly. First, most approaches require the shapes to be consistently oriented, which only applies to a very limited number of categories. Generalization behavior under unaligned training instances is much worse. Second, the correspondences induced from the point cloud generator are not good enough across the synthetic shapes, making it difficult to properly define correspondences between the dynamic template and the input fragments. Recent results in mesh representation, such as deep marching tetrahedra [197], MeshGPT [198], and MeshAnything [199], also show impressive synthesis results. However, the two limitations of the point cloud representation remain under the mesh representation. An open question is how to address these two limitations.
6 Template-Based Reassembly
We first discuss approaches that use (approximate) complete object shapes in Section 6.1. We then present approaches that leverage symmetric priors for fractured object reassembly in Section 6.2. Finally, we discuss approaches that utilize learned template shape priors in Section 6.3.
6.1 Complete object priors
In addition to feature-based reassembly, a complete object model from the same or a similar class can provide valuable prior knowledge to facilitate reassembly. For example, when reassembling objects such as skulls, pottery, and architectural elements, complete models are often available and can serve as templates to infer the initial arrangement of fragments. In recent data-driven techniques, these global geometric priors can also be learned from collected data sets using neural networks.
This section discusses how a complete object model or global prior can be used to facilitate object reassembly. We consider the input fragments as $X=\{x_{i}\}$ and the global prior as $P$. As listed in Table 2, depending on how $P$ is obtained and used, we categorize relevant methods into two types: methods using local-global alignment and methods using global evaluation. Local-global alignment methods directly use the global prior $P$ as a guide and solve the matching and alignment between $x_{i}$ and $P$ to obtain potential compositions of $\{x_{i}\}$. Global evaluation methods do not solve local-global matching, but design a function to evaluate how well a computed reassembly matches the global prior $P$. Figure 12 illustrates various methods of utilizing complete object priors for reassembly, including explicit template alignment, implicitly learned priors, and network-based global evaluation.
Methods using local-global alignment match fragmented pieces $x_{i}$ to the global prior $P$ and use this to obtain the initial or rough arrangement of the fragments. Their effectiveness depends on three factors: (1) the design of $P$, (2) the modeling/extraction of features from $\{x_{i}\}$ and $P$ for their matching, and (3) the alignment. The global prior $P$ could be a selected/computed explicit template or learned implicit parameters. Classical geometric reassembly methods often search for a good explicit shape and use it as a template or global prior, then design partial matching schemes to find local-global alignments. In contrast, as discussed in Sec. 6.3, recent learning-based methods tend to train neural networks to build implicit shape priors from datasets, and then train/optimize the encoding of both local fragments and global template so that their latent codes can infer the matching and alignment of the fragments.
In classical geometric methods, to support effective local-global alignment, the explicit shape template is often selected as an average or typical shape, or a best-matched object retrieved from a database. Using a typical object as prior shape is common in skull and cultural heritage reassembly tasks. For example, [200, 201, 189] used a standard human skull as $P$ to provide alignment guidance. [202] used a complete model from the museum’s collection as the prior to define the 3D support surface of the fragmented disk. When a larger collection of complete exemplar objects is available, retrieving a template object with the most similar geometry can provide a more adaptive global prior. For example, in the pottery reassembly field, researchers have built many digital databases of pottery. They support various types of local-global matching for the given fragments, such as drawing sketch [203], a portion of the object geometry [204], and curve descriptors [87, 205] to retrieve relevant potteries as priors $P$. In the archaeological and architectural reconstruction of a ruined monument task, to construct the most similar geometry as prior $P$, [50, 206] conducted a careful study of historical texts that mention the building under consideration. Given the selected shape prior $P$, as for the second factor, feature extraction computes the keypoints and their descriptors on both fragments and the global prior, such as depth and axial symmetry descriptor [203], salient points [204], parabolic contour [87], spin images [200, 201], Signature of Histograms of OrienTations (SHOT) descriptor [189]. With the extracted keypoints and features, fragment transformations are computed to roughly assemble the input parts to the template. Researchers developed various alignment methods using keypoint correspondences, such as iterative closed points (ICP) [200], geometric distance [201], local affine-invariant 3D moments [87], Earth mover’s distance [204], and RANSAC algorithm [189].
Implicit prior for local-global alignment is learned from an annotated 3D dataset to solve the partial matching from the fragments to the global shape. The shape prior is encoded in a neural network, which is trained on a 3D segmentation dataset, that the weights are adjusted to learn the semantic relationship between the fragments and the complete objects. In the inference stage, given a set of parts $X$, trained networks are used as encoders to encode the parts into latent codes, which inherit the implicit shape prior to the possible complete object $P$. Chaudhuri et al. [207] trained a Bayesian network on a 3D segmentation dataset to encode semantic and geometric relationships among shape components. Then a probabilistic model is used to compose and align the input parts. Mo et al. [190] proposed a more comprehensive 3D object dataset, named PartNet, a large-scale dataset of 3D objects annotated with hierarchical and instance-level 3D part information. Following the PartNet dataset, Zhan et al. [208] and Narayan et al. [209] trained an encoder on it to encode the query parts to semantic latent codes, which inherit the shape prior to the complete object. They framed the 3D part assembly problem as a graph-learning problem. They took the latent codes as input and used graph neural networks to assemble the parts into complete objects. They demonstrated their work on 3D chairs, tables, and lamps reassembly tasks.
Global evaluation as the complete object prior $P$ provides a regularization term to constrain the reassembled results. It has two variants: hand-crafted functions or learnable discriminators. Hand-crafted metrics are widely adopted in earlier reassembly tasks. In the image stitching task, an entropy-based metric is proposed to assess the quality of the stitched panoramic image [210]. In the document reassembly task, Optical Character Recognition (OCR) for word or sentence detection [211, 212]. These methods use pair-wise features to suggest candidate alignments of the fragments. They keep the candidate alignments when the evaluation results are promising, while they discard the alignments when the evaluation results are poor. Recent approaches used the capacity of large-scale datasets, trained neural networks as discriminators, and used them for global evaluation. The discriminator distinguishes the reassembled the results and the natural objects to push the network propose a better reassemble configuration that makes the results look natural. Li et al. [191] propose a GAN to solve the Jigsaw puzzles. The designed a network-based discriminator and an adversarial loss to classify the an image is reassembled or real. The adversarial loss pushes the GAN to assemble images from pieces that looks natural. Then a discriminator was trained as adversarial loss to optimize the mating configuration that should look visually realistic like an object.
Figure 13: Symmetry-based reassembly: identification and alignment using axial symmetry [213] (top) and completion based on reflectional symmetry [200] (bottom).
Figure 14: (Left) TAssembly [214]: From left to right – input, correspondence quality, assembly result, and the synthetic template shape. (Right) Jigsaw++ [215]: From left to right – input, Jigsaw++ synthetic result, and the ground truth object.
In summary, local-global alignment often works well on a specific category that has a typical complete object or a large enough database as reference. The quality of the initial alignment depends on how to encode the part-global relationship into features. Due to the lack of well-annotated database, local-global alignment might be hard to generalize to other object reassembly tasks. In contrast, the global evaluation function provides a global shape guidance to regularize the reassembled results. However, finding a proper evaluation function for efficient reassembly is still challenging.
6.2 Symmetries
6.2.1 Symmetry Detection
Axial symmetry and reflectional symmetry are the two commonly studied symmetry types in geometry analysis and processing, and especially in 3D reassembly.
As shown in Figure 13 (top), axially symmetric, also referred to as rotationally symmetric, indicates that if the complete object’s exterior surface is intersected by a plane perpendicular to its axis, the resulting cross section is a circle or closely resembles one. As a prominent geometrical property, rotational symmetry can serve as a feature to guide the 3D reassembly of fragmented relics, such as sherds. Numerous studies explored the estimation of the axis of symmetry. Earlier studies estimate the rotational axis using differential geometric properties, such as the sphere of curvature [216] and normal intersection [217]. Subsequently, enhancements of the estimated rotational axis have been proposed using M-estimators [218] or Bayesian approaches [219, 220]. However, these methods can be sensitive to geometric deformation, incompleteness, or noise, often demonstrated in fragmented pieces, leading to inaccurate detection of an object’s symmetrical axis. [221] suggested a workflow for axis estimation based on minimum thickness direction. This method does not rely on differential geometric properties, making it resilient to local defectiveness, noise, and outliers. Recent studies [222, 223, 213] have broadly adopted the random sample consensus algorithm (RANSAC) to find good initial estimates. [222] extended the Cao and Mumford’s method [216], to consider both the inner and outer surfaces. This technique introduces PotSAC, a two-stage axis estimator based on a variant of the RANSAC algorithm followed by a robust nonlinear least-square refinement, making it capable of handling noisy surface normals. [223] applied beam search to explore multitudes of registration possibilities, addressing false matches between sherds inflicted by indistinctive sharp fracture surfaces. [213] showed that the identification of the symmetry axis from surface normal lines may yield two meaningful local minima. A multiple-solutions RANSAC algorithm is proposed for initial estimates, followed by a coordinate-descent algorithm to refine these estimates. The authors showed that their method ensures convergence, achieves a faster convergence rate, and provides state-of-the-art performance in axis estimation.
As illustrated in Figure 13 (bottom), reflectional symmetry refers to the property that an object does not change upon undergoing a reflection. Human skull typically exhibits reflectional symmetry along the mid-sagittal plane [224]. This kind of symmetry is one of the main geometric priors used in skull reassembly and completion [201, 200].
6.2.2 Symmetry-guided reassembly
With an estimated symmetry axis, the consistency of the symmetry axis and the axis profile curve can be viewed as a constraint or optimization objective during reassembling. [225, 226] implemented a global constraint that all fragments share the same axis, greatly reducing computational costs and improving noise resistance. This method reduces the search space from 6 DOF (for two fragments) to 1 DOF (relative position of two pieces along the axis of symmetry). [223] relaxed the above constraint by encoding the information about the axis as a descriptor into each break line for efficient reassembly.
6.3 Learned Shape Spaces
Compared to approaches that rely on similar shapes or symmetric templates to guide fractured object reassembly, few methods have explored the use of a generic template shape space learned from the data to enhance this process. One notable exception is TAssembly [214], which learns a linear shape space specifically for bones. It formulates fractured object reassembly as an optimization problem, adjusting fragment poses and the underlying complete shape in the learned shape space to minimize the distance between the fragments and the dynamic template. However, this method is limited to bone shape spaces, where consistent mesh parameterization makes it relatively straightforward to learn descriptors for predicting dense correspondences. Another promising approach is Jigsaw++ [215], which employs a point cloud generation model with a rectified flow-based retargeting algorithm to infer the complete shape from a partially assembled input. Although this approach shows significant potential for reconstructing daily objects, scaling it to handle more diverse or even open-vocabulary input shapes requires larger models and more data. The challenge remains to generate shape priors when the underlying shape space exhibits large geometric and topological variations. Another limitation is that it does not compute correspondences between the reconstruction and the input fragments. Therefore, it does not enforce the rigidity constraints among fragments.
7 Applications
We discuss five major application areas of fractured object reassembly (Section 7.1 to Section 7.5) and the relevant application in shape modeling via part assembly.
7.1 Bone Reduction
3D bone reduction serves as a key step in preoperative surgical planning [227, 228, 229] and constitutes a critical element of intraoperative navigation to reduce fractures [230, 231] or a robotic system [232, 233](see Figure 15). The objective of bone reduction is to restore the native anatomical configuration of bony structures, encompassing both individual bones and articulations. Unlike other fractured entities, fractured bones exhibit the following characteristics. Primarily, human bones exhibit pronounced shape priors, which can be leveraged to facilitate the reassembly procedure (see Section 6). Owing to the inherent reflection symmetry of human anatomy, the intact contralateral bones can serve as the archetypal reference for the pre-fracture bone structure. Secondly, unlike other objects, bone is composed of compact (cortical) bone, cancellous tissue, and cartilage. In instances where the fracture occurs within the cancellous tissue, especially within the metaphyseal region of a long bone, the resultant fracture interface commonly presents considerable indeterminacy, which affects the precise alignment of bone fragments [214, 234]. In the literature, the general workflow of bone reduction includes four components, i.e., CT bone segmentation, template matching, pairwise fragment matching, and multi-fragment matching.
Figure 15: The surgical robot designed for bone reduction is detailed in [235, 236]. At the top, the figure illustrates the components of the surgical robot during a bone reduction procedure within a surgical setting. At the bottom, the figure presents the surgical planning system associated with the robot.
CT bone segmentation endeavors to delineate the target fractured bone from adjacent healthy bone and to discriminate the bone fragments from each other in CT images. Subsequent to delineation, bone fragments are commonly divided into the fracture zone and the intact zone, a process that constitutes one of the most arduous challenges in CT bone segmentation. The methods identifying fracture zone are comprised of automatic approaches such as watershed [227], probabilistic watershed [237], template-matching [234] or semi-automatic methods such as region growing [238, 239] or continuous max-flow [230]. In addition to the non-DL methods mentioned above,  [240] implements supervised contrastive learning with a novel Siamese deep neural network so as to identify and localize the pelvic fracture zones. In [234], an approach was introduced to synergistically segment and reassemble the fragments utilizing the contralateral bone as a template and exploiting the synergistic interaction between segmentation and reduction (see Figure 16). In recent advancements delineated in [241], an innovative dual-stream learning framework has been introduced for the automated segmentation and categorical labeling of pelvic fractures. This framework facilitates the identification of pelvic fracture fragments in varied quantities and anatomical locations through a dual-branch architecture that harnesses distance learning based on bone fragments. Considering that the bone template is constituted by the outer surface of the intact zone, aligning the bone fragment with the template necessitates the segmentation of the outer surface of the intact zone [237, 234]. The bone surface can be extracted by the MarchingCube algorithm [47].
Template matching is employed as an initial coarse reduction step by aligning the bone fragments with the bone template. The healthy contralateral mirrored bone frequently serves as a reference template for the reduction of the fractured side [227, 234, 242, 237, 243, 244]. Matching the bone fragments with the contralateral bone template can be formulated as a part-to-whole matching problem for 3D rigid objects. However, the existence of intra-individual variations, particularly discrepancies in axial torsion, led [245] to conclude that preoperative planning methodologies aimed at reconstructing intricate proximal humerus fractures should not depend uncritically on contralateral anatomical structures. Furthermore, contralateral bone is not always available due to ethical considerations in medical practice and instances where the pathology affects both sides. Due to the similar morphological characteristics of human bones, the parametric statistical shape model (SSM) serves as an alternative template for fracture reduction [214, 228, 230]. The alignment of bone fragments with the parametric morphable template is formulated as a part-to-whole problem for 3D non-rigid objects, whereby the output encompasses the rigid 6D pose of each bone fragment and the deformation parameters of the SSM model. Moreover, to address the complexities associated with joint dislocation, a statistical pose model (SPM) has been meticulously developed to accurately represent the anatomical pose of the bones comprising the bony joint [246]. Significantly, the seminal work by [230] integrates both the SSM and the SPM frameworks to address intra-bone fractures and inter-bone dislocations. Given the inherently non-convex nature of this complex problem, the iterative procedure alternates through three critical stages: fragment alignment update, SSM adaptation, and SPM adaptation.
Figure 16: Synergistically segment and reassemble the fragments by utilizing the contralateral bone as a template [234].
Pair-wise matching aims at aligning two fragments so that their fracture boundaries can perfectly match. Pair-wise matching can be divided into two steps: identification of fracture boundary and matching of the fracture boundary. After template matching, the fracture boundaries between fragments are close. Subsequently, the Euclidean distance between the surface of two fragments can be used as a criterion to identify the fracture boundary [227]. In instances where fractures occur within rigid regions such as cortical bones, the resulting fracture boundary typically manifests pronounced curvature. Subsequent studies [242, 247, 248, 249, 237, 250] have used this characteristic to effectively identify fracture boundaries. In contrast, in cases where fractures occur within more flexible regions, such as trabecular bone, the fracture surface tends to appear relatively planar, making the fracture boundary less distinct. In such circumstances, interactive annotation of the fracture boundary becomes indispensable [251, 242]. In the context of matching fracture boundaries, if the fracture occurs within the rigid cortical bone, the methodologies for pair-wise matching bear significant resemblance to those employed in the reassembly of precise artifacts, such as cultural relics. These techniques [83, 238, 249] frequently use criteria in which tangential contact points on the respective bone fragments converge at the coordinates of the coincident point, with surface normals oriented oppositely, and matching surface curvatures exhibiting inverse signs. Concurrently, the Iterative Closest Point (ICP) algorithm is an effective instrument to execute pairwise matching [252, 242, 248, 250, 227]. In contrast, in instances where the fracture occurs within the more pliable, trabecular bone, or in scenarios where diminutive fragments are absent near the boundary, pair-wise matching is likely to be ineffective. In such circumstances, the previously discussed template matching emerges as an essential and indispensable methodology for bone reduction. Note that while common pairwise registrations and pairwise matching techniques such as ICP and feature matching are used in pairwise bone fragment matching, the approaches introduced above instill domain-specific knowledge of individual bones.
Multi-fragment matching, recast as a quintessential NP-hard problem, seeks to globally optimize the poses of multiple fragments. The input for multi-fragment matching includes the current poses of all bone fragments following template matching and pair-wise matching. The objective of global optimization is to maximize the region matching score, which combines the scores of template matching (if a template exists) and pair-wise matching [189]. To achieve this, existing methodologies typically employ an exhaustive search [238] or a greedy search strategy [253] incorporating heuristic rules to address the aforementioned global optimization. Within the greedy search strategy, bone fragments from the inactive set are sequentially added to the expanding set of already aligned objects, while ensuring the correct alignment of objects within the active set. The next fragment in the inactive set is selected based on the highest score, subject to penetration validity checks. In [189], the score was calculated as the weighted sum of the scores from prior template matching and pair-wise matching. Furthermore, metrics such as the distance between the fragment and the active set [247], the area of the fracture surface [250], and the size of the bone fragment [227] were also used as priority scoring criteria. To avoid entrapment in local minima, multiple fragments can be chosen, rather than a single fragment, to add to the active set using the beam search approach [189]. A notable challenge in multi-fragment matching is the potential accumulation of pairwise/template errors in multi-piece alignment. To address this issue,  [247, 237] introduced the ideal mates concept, which enables the calculation of the multi-piece alignment error without depending on the pairwise registration error [253]. Unlike [247, 237][189] implemented graph optimization on the target graph, representing the final correspondences among the assembled bone fragments, to globally refine existing transformations. Similar to pairwise matching, there are similarities between the multi-fragment matching approaches introduced for bone reduction and those used in restoring stone fragments. However, most approaches to bone reduction consider domain-specific knowledge.
Deep learning in bone reduction. In the literature, the deep learning models were trained to segment the cortical surface of bone fragments [234] and fractured bones [254], identify and localize bone fractures [240], and compute bone features [214, 234] for template matching. Recently, a significant study [231] introduced an innovative bidirectional framework for automated pelvic fracture surgical planning, which integrates fracture simulation with structural restoration. The fracture simulation module specifically considered patient-specific pelvic anatomy, bone density variations, and fracture stochasticity to generate diverse fracture scenarios from intact pelvises. For the creation of training datasets that incorporate actual fractures, images of the fractured side and contralateral healthy side were utilized as input and ground-truth data, respectively. Accordingly, being trained on both simulated and real fractured bones, the proposed deep structural restoration network reconstructs the bone’s pre-fracture state, subsequently serving as the template for matching the input bone fragments.
7.2 Fractured Thin-shell Object Reassembly
In the context of 3D reassembly tasks, a thin-shell object refers to a structure with a thickness that is negligible compared to its other dimensions. This implies that the object’s surface geometry, rather than its volume, primarily defines its shape. Thin-shell objects present significant challenges during reassembly due to their intricate nature, making the matching and alignment of fragments more complex.
In forensic skull reassembly, the skull fragments are considered thin shell objects because their thickness is minimal compared to their overall size. This requires a precise matching of the fracture regions and intact surfaces to accurately reconstruct the original shape. Yin et al. [200] and Wei et al. [201] model the set of skull fragments as a collection of surfaces. They detect the boundary loops of the skull fragments as feature curves and then compute feature descriptors on these curves. The feature curves are used for pair-wise matching and alignment computation. However, Zhang et al. [189] note that due to the thickness of the skull fragments, the feature curves can form multiple loops and may be ambiguous in a single fragment. To match multiple feature curves, they formulate it as the Largest Common Point-set (LCP) problem. They developed a voting algorithm using 4-point congruent sets extraction to solve this LCP problem.
Figure 17: Photo from [87] that shows a project that reconstruct thin-shell objects from its fragments. (Left) fragments. (Right) A digitized fragment.
Another relevant scenario is archaeological artifact reassembly such as pottery [87, 255], where broken fragments are often thin. Cohen et al. [87] first digitized the excavated fragments using a Konica 3D scanner (See Figure 17). They treated the scanned fragments as surfaces and computed the parabolic contours on them. Finally, they used affine-invariant 3D moments to match the contours and align the fragments. Their system shows a magnitude improvement in running time compared to manual solutions.
Angelo et al. [256] proposed a recognition pipeline to determine the regions of the 3D scanned fragments into the rim, internal/external wall, and base. To reassemble a complete pottery, they designed various descriptors for different regions of the fragments. These descriptors help find correspondences and align the fragments. Fu et al. [257] take curvature and torsion as characteristic descriptors of the curve and search for the candidate matching by the derivative dynamic time warping (DDTW). They then design a voting mechanism and ICP method to align the fragments. Wang et al. [24] observed that due to the thin shell nature of pottery fragments, thin side strip regions often contain many repetitive fractured patterns and lack sufficient unique geometric features. They provide a novel contour-based solution to solve batch matching and a boundary-aware ICP method to tackle fine registration. They devised an intrinsic shape descriptor to encode a 2D contour to fuse the front and back scans. After each fragment’s front and back scans are paired and initially aligned, to address the issue of small overlap and the lack of salient textures and features in the small overlapping region of the two scans, they propose a boundary-based ICP method to ensure robust ICP iteration. Their approach demonstrated the ability to scan and reconstruct over 700 fragments in 10 working hours, achieving high reconstruction accuracy.
7.3 Wall Fragment Reassembly
Figure 18: The top row shows projects of restoring wall fragments. The bottom row shows projects of restoring building fragments. (Top-left). Stanford’s Forma Urbis Romae project [258]. (Top-right) Princeton’s . Theran wall paintings project [22]. (Bottom-left). Restoration of the Octagon monument in Ephesos [50]. (Bottom-right) Reconstruction of the Arch of Titus at the Circus Maximus in Rome [206].
A special class of thin-shell fragments are wall fragments, in which the complete objects is flat. The first well-known effort is Stanford’s Forma Urbis Romae project [258] (Figure 18 (Top-Left)). They scanned 1186 fragments that are physically available and, according to estimates, account for about 20% of all fragments. As a significant portion of the fragments are missing, they only report dozens of potential matches [79]. The methods to compute use methods include boundary incision matching, which employ surface textual features, wall feature matching, which is based on location prediction using a underlying map, and multivariate clustering. Multivariate clustering finds fragment pairs that may not have explicitly corresponding incisions, but may share other common characteristics including fragment thickness, marble veining direction, axial direction of the incised architecture, presence and orientation of slab edges, and the back surface condition of the fragments (rough, smooth, or unpreserved.
Another category of notable projects are efforts to reconstruct and match Theran wall paintings [259, 260, 22, 261, 140, 262, 144], including a series of work at Princeton University on reconstruction [22], matching [260, 261, 140], and analysis of fragment boundary patterns for fragment pattern simulation [262]. Moreover, [144] presents a generic algorithm for reassembling wall paintings. One major difference between wall fragments and fragments of bones and thin-shells is that the number of wall fragments is much larger. In other words, it is important to develop computational solutions to fractured object reassembly. Another consequence of having a large number of fragments is that the global consistency constraint [140, 144] becomes important for multi-fragment matching.
One relevant scenario is the restoration of ancient papyrus fragments, which share the common characteristics of being flat. [263] addresses the reassembly of these fragments using a convolutional neural network in conjunction with a random forest classifier.
7.4 Fractured Building Reassembly
Another special class of fragments are fragments of architectural buildings. In this case, virtual restoration is necessary as it is very difficult to move the fragments. Similar to the setting of wall fragments, the number of fragments is much larger than in the case of bone fragments or thin shell fragments. Moreover, the percentages of missing fragments in the building setting, due to the open environment, are even higher than those in the wall painting setting. In this case, solving the fractured object reassembly problem strongly rely prior knowledge of the underlying complete object. Moreover, user input, i.e., semi-automatic approaches, are necessary to obtain good results.
One example is the project to reconstruct the Octagon monument in Ephesos [50]. The project used a semi-automatic approach to virtually restore the original monument from 130 fragments (about 20% of the total number of fragments). This restoration is heavily based on prior knowledge of the monument. Another relevant project is the reconstruction of the Arch of Titus at the Circus Maximus in Rome [206]. Again, many fragments are missing, and the virtual reconstruction relies on knowledge about the underlying symmetric structure of the Arch. Other notable efforts include the restorations of mixtilinear buildings in Villa Adriana [264], Macedonian Tombs [265], the Maritime Theatre at Hadrian’s Villa [266], and a project on restoring damaged column drums  [267].
7.5 Fractured Object Reassembly in Science
There are many scientific applications for fractured object reassembly. One such example is docking of proteins and DNA Sequencing. In [268], in which mitochondrial genes from the unicellular eukaryote Diplonema are systematically fragmented into small pieces that are encoded on separate chromosomes and individually transcribed. The reassembly problem is to concatenate them into contiguous messenger RNA molecules. Another biological application is protein docking [269, 270], where the goal is to find a rigid and non-rigid transformation of one 3D protein to interact with another protein.
Another application is in paleontology, in which we can reconstruct fossils from fragments precisely in 3D, to allow 3D analysis of the whole shape. A semi-automatic approach was introduced in [271] to reconstruct an extinct bird. It is desirable to develop automatic approaches for more accurate and faster reconstructions. [272] studied virtual reconstruction of huminin crania to understand human evolution. [273] introduced an interactive approach in Blender to restore a dorsal rib and frontal bone of a small-bodied Jurassic-age armored dinosaur from Africa, The digital restoration engaged all modalities of deformation (translation, rotation, scaling, distortion) and reconstruction (fracture infilling, adding missing bone, surface smoothing).
Figure 19: Applications in scientific domains. (Top-Left) DNA sequencing [268] (Top-Right) Protein docking [270] (Bottom-Left) Fossil reconstruction of a bird [271]. (Bottom-Right) Volcano crystal reconstruction [274].
One other potential scientific application is the reconstruction of crystals that “blew up” during eruption because of the fast decompression of trapped gases [274]. In this setting, it is important to reconstruct not only the particles but also the voids within them that originally held the gases. Such reconstructions have great implications in geosciences.
7.6 Part Assembly
Part assembly refers to the process of assembling components that are decomposed in a semantically meaningful way, such as furniture components like table legs[190] and mechanical parts like screws[275]. Although closely related to the fracture assembly problem, part assembly introduces several key differences. It generally assumes that each part has an independent semantic function and that each component is complete and functional on its own, unlike the fractured pieces in fracture assembly. Moreover, part assembly allows for duplicated components, and solutions are often non-unique. Unlike fracture assembly, where local or global geometry provides significant clues, part assembly relies more on semantic or functional understanding of the individual parts and the assembled object. However, despite these differences, we can identify valuable connections between the two problems, and both fields can benefit from exploring the shared foundations of these challenges.
A notable branch of part assembly research is the object assembly problem, which focuses on assembling parts to create complete daily objects such as tables, lamps, and chairs. Early work in this domain addressed two primary questions: how to collect relevant parts and how to combine them effectively. The foundational work in this area was proposed in Modeling by Example [276], where the authors introduced a method to retrieve parts from a repository of segmented 3D objects and combine them to create new objects from different parts. Building on this, SnapPaste [277] further developed the concept by focusing on how to align and model the mesh at the connecting areas of these parts. Several subsequent papers [278, 278, 279] introduced graphical probabilistic models to improve the selection and organization of retrieved parts. Additionally, research such as [280, 281, 282] explored retrieving or generating parts in an assembly process to complete objects from partial input, image data, or artist-created models. These developments laid the groundwork for later efforts to apply object assembly techniques to 3D asset generation [283, 284, 285].
Figure 20: Part Assembly examples. (Top) Object assembly [286]. (Bottom) Mechanical assembly [275].
More recent work in object reassembly has shifted towards estimating the 6-DoF pose of input parts. Unlike earlier methods, these approaches typically assume that the provided parts are sufficient to form a complete object, similar to the process of assembling IKEA furniture. In support of this area, IKEA has released the IKEA 3D Assembly Dataset [287], which includes models and assembly manuals for six of their products. Additionally, the IKEA ASM Dataset [288] focuses on assembling IKEA objects using human demonstrations. On a broader scale, PartNet [190] has become a major benchmark for this research area. For example, Li et al.[289] used 2D images to guide the estimation of 3D poses for input parts. Techniques such as dynamic graph learning (DGL) [208] were proposed to generate part pose proposals, with later works incorporating RNNs [209] and transformers [290, 291] for improved accuracy. Researchers have also refined these methods by incorporating matching information [292] and hierarchical structures [286] to improve assembly precision.
Another key focus is the mechanical assembly problem, which differs from object assembly in that the provided parts are complete functional units, and the goal is to build a working mechanism. This problem often involves parts in the B-Rep format specific to mechanical engineering. Although this is a relatively new field, several papers have addressed different aspects of the problem. AutoMate [293] presented an assembly-manual-based dataset of mechanical assemblies with 92,529 instances and proposed a graph convolutional network to predict mating locations and types for two input parts. JoinABLe [294] introduced an assembly dataset from the Fusion360GalleryDataset with 8,251 assemblies and 154,468 separate parts, along with a joint axis prediction network and a joint pose search algorithm based on predicted axes. Assemble Them All [275], which included 12,970 assemblies, addressed the challenge of multi-part, multi-scale input data with a disassembly-guided BFS algorithm.
Figure 21: Datasets that are suitable for deep learning research in fractured object reassembly. (Top-Left) The real Fantastic Breaks dataset [295] which combines both isolated fragments (top) and paired complete objects and their fragments (bottom). (Top-Right) A real dataset of fragments with ceramic material [24] is captured using the system shown here. (Bottom) A synthetic dataset created by performing fracture simulation on a diverse set of objects [145].
Apart from these two areas, object assembly also emerged in LEGO assembly. [296] proposed a segment and pose estimation network to translate the LEGO manual to executable planing, and StableLego [297] introduced an evaluation method to assess the stability of a LEGO assembly.
Implicit in these assembly approaches are priors about the underlying complete objects and priors on how pairs or a subset of primitives interact. These two pairs are related to the global template shape priors and pairwise fragment matching in fractured object reassembly. As stated previously, the major difference is that part assembly focuses on semantic while fracture reassembly focuses on geometry.
8 Datasets and Software Packages
We first discuss existing datasets in Section 8.1. We then discuss existing software packages 8.2.
8.1 Datasets
Machine learning approaches have shown great advancements in the task of reassembly of fragments, but are highly dependent on large-scale datasets. However, the first data sets used in fragment reassembly research are primarily small-scale, often manually created [52, 298, 258] or derived from automated semantic segmentation [293, 190].
Figure 22: Illustrations of future directions (Left) Representations and algorithms for matching complete template shapes with large geometric and topological variations and end-to-end learning of the entire reassembly pipeline. (Middle) Leveraging foundation models to solve sub-tasks such as segmenting fractured surfaces, matching fragments, predicting complete template shapes from input fragments, and matching complete template shapes and fragments. (Right) Learning to generate synthetic datasets of 3D fragments and and doing so jointly with learning the reassembly network.
There are two recent efforts which have focused on scanning fragmented objects at scale. The first example is Fantastic breaks [295], which include a small collection of objects with broken fragments and more objects that were broken into fragments by the authors (see Figure 21(Left)). Both fragments and complete objects (if they exist) are scanned. Another example is [24], which developed a system to batch scan fragments of objects with ceramic material (see Figure 21(Right)). This dataset has ground-truth complete objects for evaluation. In addition, the scale of these two datasets makes it possible to develop deep neural networks for reassembly. However, the number of categories in these two datasets is still small, presenting generalization issues in novel categories.
In addition to real data sets, Breaking Bad [145] is a synthetic data developed by running a simple fracture simulation procedure to generate fragments of shape collections that undergo large geometric and topological variations ((see Figure 21(Right)). This dataset is interesting for fractured object reassembly because the underlying objects belong to diverse object categories. For example, template-based reassembling techniques can be developed by learning a shape prior of the complete objects in this dataset. Just as the role of ShapeNet [299] in boosting 3D deep learning rich, we envision that research progress on algorithmic fractured object reassembly will be accelerated with Breaking Bad and similar datasets. However, as the fragments are generated using non-physics based fracture simulation. The question is how the reassembly pipelines derived from such synthetic datasets generalize to real data.
8.2 Software Packages
Few early approaches to fractured object reassembly released code. A major factor is that in the early days we did not have established benchmark datasets for experimental evaluation. For example, there are rich benchmark datasets in computer vision and people deliver open-source codes that execute on benchmark datasets. With the introduction of large-scale benchmark datasets, we see the trend to release codes. For example, many papers that evaluate using Breaking Bad have open-source implementations, including Jigsaw [61], SE(3) equivariance [142], DiffAssemble [153], and PuzzleFusion++ [141].
9 Future Directions
We are in the deep learning era, in which there are breakthrough results after breakthrough results on developing deep learning approaches to solve core problems in geometry processing and 3D vision. Undoubtedly, the future of tackling the reassembly problem lies in deep learning approaches. Although we have discussed many recent advances on this front, there are open questions from neural representations to algorithms to training data, which should be addressed before deploying reassembly pipelines in practice.
On the representation and algorithm side, there are open questions on how to learn shape priors of diverse complete objects. When targeting general object categories that have large topological and geometrical variations, implicit shape representations, which can encode shapes with varying topology, become the only choice. However, it is an open problem to properly define correspondences between fragments, which are normally represented as point clouds.
Once we have addressed this open problem, it makes sense to study how to integrate all components of the reassembly pipeline in an end-to-end manner, enabling effective training. Currently, it remains an open question on how to do so when the reassembly pipeline uses deforming template shapes.
In addition, having sufficient training data is critical to the success of developing deep neural networks. However, this becomes challenging in the case of the fragment reassembly problem. First of all, the number of complete objects that have fragments remain sparse. In addition, developing software and hardware systems to scan these fragments is very costly. There are two interesting topics related to addressing this data issue.
The first topic is on how to boost the performance of reassembly using vision foundation models. Many sub-tasks in the reassembly pipeline, including fracture surface segmentation, establishing correspondences between pairs of fragments, predicting the underlying complete object from fragments, and matching fragments with a complete object, have counterparts as core geometry processing and 3D vision problems. Recently, we have seen that 2D foundation models, either through rendered images or lifted 2D features, have led to breakthrough results on these tasks. This brings interesting questions on how to adopt these approaches on the corresponding sub-tasks in reassembly. For example, a recent paper [300] has shown that vision-language models can effectively understand the material properties of objects from rendered images. Note that rendering images of fragments is fundamentally different from rendering images of complete objects for visual understanding. How to do so to reveal fracture regions is a very interesting research question.
The second topic is on generating synthetic fracture data to train fracture reassembly networks. The advantage of this approach is that we have abundant complete 3D objects, allowing us to learn shape priors. Moreover, we have ground-truth poses and correspondences from synthetic data, which are hard to obtain from real data. In fact, several synthetic datasets [145] have been introduced in the past few years. As discussed in this survey, these datasets have led to a wave of research activities on developing deep learning pipelines to solve the reassembly problem. However, a fundamental challenge when applying these deep reassembly pipelines in practice is the domain gap between synthetic data and real data. To this end, it is important to develop an end-to-end learnable fracture simulation pipeline, so that we can control the characteristics of the simulation output, e.g., the number of fragments and erosion among fracture surfaces. This network should be physics-aware, so that we do not need to generate a lot of accurate but costly training data for supervision. This network should also properly model the weathering effects, which are prevalent in real datasets. Such a network allows us to learn hyper-parameters of fracture simulation, e.g., material properties and distribution of external forces, from simulated data to maximize its performance on real data.
Once we have an end-to-end learnable fracture simulation pipeline, another important problem is to study how to combine fracture reassembly and fracture simulation in an end-to-end manner. This allows us to optimize the fracture simulation network, so that the fracture reassembly network generalizes optimally when it is trained from the fracture simulation output. This task goes beyond the standard notion of end-to-end training, as we have to develop losses on the output distribution of the fracture simulation network.
From a community-building perspective, it is important to create large-scale benchmark datasets for experimental evaluation. This will attract more people to work on this problem and, according to the trend in computer vision, will simulate more people to release open-source codes. We do observe a good initial trend from the Breaking Bad dataset [145].
Finally, it is also important to identify more applications of fractured object reassembly in scientific domains. This will amplify the broader impacts of this research community.
10 Conclusions
In this survey, we have provided a thorough survey on the reassembly problem, which finds rich applications in cultural heritage restoration, bone fracture restoration, and object/scene modeling from parts and objects. We have covered early non-deep learning based approaches to more recent deep learning based approaches. Due to connections to core geometry processing problems in shape segmentation, shape matching, symmetry and primitive detection, and learning shape priors, we also discussed relevant approaches on these problems, emphasizing similarities and differences. We also covered current state-of-the-art methods for solving real problems in these application domains.
We envision that this topic will undergo significant progress in the coming years from the perspective of developing deep learning based approaches. We have highlighted two promising venues for future research. The first studies how to leverage foundation models to boost various sub-problems of reassembly. The second examines how to generate synthetic data in a controlled manner, which enables joint learning of reassembly and fracture simulation. We conclude that this very fundamental research field can also benefit from having more scientific applications.
Acknowledgement. Qixing Huang is supported by NSF Career IIS-2047677, NSF IIS-2413161, and Gifts from Adobe and Google. Xin Li is partly supported by NSF CBET-2115405 and NIH R15HD108765. Junfeng Jiang would like to acknowledge support from Jiangsu Province Key Research Program-Social Development-Clinical Frontier Technologies (BE2022718).

### v2

@@
 The purpose of this article is to provide a systematic survey on this topic.
-The reassembly problem requires understanding the attributes of individual pieces and establishing matches between different pieces.
+This reassembly problem requires understanding the attributes of individual pieces and establishing matches between different pieces.
 Many approaches also model priors of the underlying complete object.
@@
 We provide existing algorithms in this context and emphasize their similarities and differences to general-purpose approaches.
-We also survey the trends from early non-deep learning approaches to more recent deep learning approaches.
+We also survey the trends from early procedural approaches to more recent deep learning approaches.
 In addition to algorithms, this survey will also describe existing datasets, open-source software packages, and applications.
@@
 Reconstructing a complete object from its parts is a fundamental problem in many scientific domains.
-In archaeology, we typically collect fragments of one or multiple objects at an archaeological site, and the goal is to restore the original objects.
-When people have bone fractures, we need to study how to assemble bone fragments to plan bone reduction surgery.
-This problem also arises in the 3D object/scene modeling, where we want to construct a complete object/scene from parts/objects.
-Other relevant problems are protein docking and human object interactions, in which the difference is that the transformation of each part becomes non-rigid.
-The purpose of this article is to provide a systematic survey on this topic, including algorithms, datasets, open source software packages, and applications.
+In archaeology, we collect fragments of one or multiple objects at an archaeological site, and the goal is to restore the original objects.
+In medical applications, we need to study how to assemble bone fragments to plan bone reduction surgery.
+This problem is also crucial in paleontology for reconstructing fossils and in forensics for evidence reconstruction.
+The challenge extends to 3D object/scene modeling, where we want to construct a complete object/scene from parts/objects.
+Other relevant problems are protein docking and human-object interactions, in which the difference is that the transformation of each part becomes non-rigid.
+The purpose of this article is to provide a systematic survey on this topic, including algorithms, datasets, open source packages, and applications.
 To the best of our knowledge, this is the first systematic survey on this topic.
 The reassembly problem is related to many shape analysis tasks.
-An important task is to segment the regions of the fractured surfaces and the original surfaces, which is a special shape segmentation problem.
+An important task is to segment and classify the parts’ surface into fractured and intact regions, which is a special shape segmentation problem.
 Another important task is to find correspondences between fractured regions of different pieces to find relative poses between them.
-This is related to scan matching, with the difference in that the correspondences are between complementary points.
+This is related to scan matching, with the difference that the correspondences are between complementary points.
 In addition, many approaches utilize priors of the underlying complete object, including symmetries and shape priors.
 This calls for approaches to analyze and encode such priors and use them to solve the reassembly problem.
-Similar to other problems in computer graphics, we have also witnessed the transitions from early non-learning based approaches to recent deep learning approaches.
-Early approaches have explored all aspects of the reassembly problem, including segmentation of fractured regions and original regions, matching fractured regions, and encoding priors (symmetries and shape spaces) of the underlying complete shape.
-However, they are restricted to hand-crafted features, so the entire systems do not generalize robustly.
+Similar to other problems in computer graphics, we have also witnessed the transition from early procedural approaches to recent deep learning approaches.
+Early approaches have explored all aspects of the reassembly problem, including the segmentation of fractured regions and original regions, matching of fractured regions, and encoding of priors (symmetries and shape spaces) of the underlying complete shape.
+However, they are restricted to hand-crafted features, so entire systems do not generalize robustly.
 In addition, it is difficult to encode shape priors among objects with large geometrical and topological variations.
 These issues can be addressed using deep learning approaches.
-Besides the fact that the shape segmentation and shape matching problems become more robust, we also have approaches that perform shape segmentation and shape matching together, which benefit from end-to-end training in deep learning.
-In addition, with advances in learning 3D shape generative models, it is possible to learn shape priors from diverse shape collections to reassembly.
-Figure 1: Overview of this survey. (a) Single-piece processing and analysis. (b) Multi-piece analysis. (c) Template shape space and template-based reassembly. (d) Applications.
+In addition to the fact that the shape segmentation and shape matching problems become more robust, we also have approaches that perform shape segmentation and shape matching together, which benefit from end-to-end training in deep learning.
+In addition, with advances in learning 3D shape generative models, it becomes possible to apply learned diverse shapes prior to reassembly.
+Figure 1: Overview of the methods presented in this survey. (a) Single-piece processing and analysis. (b) Multi-piece analysis. (c) Template shape space and template-based reassembly.
 Despite tight connections to relevant problems in shape analysis, applying existing approaches to the 3D reassembly problem is not straightforward.
-This is the case for both non-deep learning and deep learning based approaches.
-However, in the literature, people do not establish the reassembly as a unique sub-area in geometry processing and shape analysis.
-Our goal is to summarize state-of-the-art results in this domain systematically, seeking to establish the skeleton of this important sub-area.
+This is the case for both procedural and deep learning based approaches.
+While reassembly shares many technical components with other geometry processing tasks, it has not been formally recognized as a distinct sub-area within geometry processing and shape analysis.
+Through this survey, we aim to systematically organize state-of-the-art results in this domain and establish a structured foundation for this important research direction.
 1.1 Related Surveys
@@
 The difference between our survey and [1] is two-fold.
-First, in contrast to their survey which focuses mostly on non-deep learning techniques, our survey covers recent advances using deep learning.
+First, in contrast to their survey which focuses mostly on procedural techniques, our survey covers recent advances using deep learning.
 Second, in contrast to the specific application of assembling fractured objects, this survey covers many applications, including bone fracture reduction, shape modeling by assembling parts, and applications in biology, geoscience, and paleontology.
@@
 In contrast, we focus on solving the inverse problem of reconstructing complete objects from parts.
+[3] discussed techniques for reconstructing fractured archaeological textiles.
+It focuses more on using textiles as guidance for puzzle solving and archaeological reassembling in both 2D and 3D cases.
+Our work demonstrates more general cases and applications where textiles might not exist.
 Our survey also discusses geometry processing techniques, including 3D reconstruction [4, 5], mesh segmentation [6, 7], primitive segmentation [8], shape descriptors [9] and correspondences [10], symmetry detection [11], and data-driven shape analysis and processing [12].
-While the pipeline of reassembling parts uses these techniques, this article focuses on approaches tailored for this particular class of applications, which usually have modified inputs, outputs, and problem setups.
+Although the process of reassembling parts uses these techniques, this article focuses on approaches tailored for this particular class of applications, which usually have modified inputs, outputs, and problem setups.
 Many reassembly techniques utilize shape priors of the underlying complete priors.
@@
 There are also other relevant review articles.
-[15] surveys related approaches for each step in the pipeline that starts from reassembly to object completion, which also addresses missing regions.
+[16] survey related approaches for each step in the pipeline that starts from reassembly to object completion, which also addresses missing regions.
 [17] present a survey article on the specific problem of reassembly of pottery fragments.
-[17] surveys related work on reassembly of defective cultural heritage objects.
+[18] survey related work on reassembly of defective cultural heritage objects.
 Our survey presents a comprehensive review of general fractured object reassembly and connections to relevant research results in geometry processing.
 1.2 Paper Organization
-We organize the reminder of this paper as follows.
+We organize the remainder of this paper as follows.
 In Section 2, we provide a more detailed overview of this paper.
@@
 In Section 7, we discuss various applications of the reassembly problem.
-Section 8 present available datasets and software packages.
+Section 8 presents available datasets and software packages.
 In Section 9, we present current trends and future directions on this topic.
 Finally, we conclude this survey in Section 10.
+Figure 2: Overview of applications presented in this survey.
 2 Overview
-This section presents an overview of this survey (See Figure 1).
+This section presents an overview of this survey (See Figure 1 and Figure 2).
 Single piece processing and analysis (Sec.
 3).
-Single piece processing and analysis concerns individual pieces of fractured objects.
+Single piece processing and analysis concern individual pieces of fractured objects.
 Topics include how to digitize physical fragments into digital 3D models and how to segment each fragment into regions of original surfaces and fractured surfaces.
-Although these topics are central topics of the field of geometry processing and shape analysis, we focus on techniques that are tailored for the particular characteristics of fragments of physical objects.
-We will cover both non-deep learning approaches and more recent deep learning approaches.
+Although these topics are central topics in the field of geometry processing and shape analysis, we focus on techniques that are tailored to the particular characteristics of fragments of physical objects.
+We will cover both procedural approaches and more recent deep learning approaches.
 Multi-piece analysis (Sec.
@@
 Multi-piece analysis studies the critical problem of establishing correspondences between fragments to match them.
-We begin with pairwise matching techniques using geometric and textual features extracted from the fracture surfaces and original surfaces.
+We begin with pairwise matching techniques using geometric and textural features extracted from the fracture surfaces and original surfaces.
 We will survey global techniques that do not require prior knowledge of the relative pose between a pair of fragments and local techniques that start from reasonable initial relative poses.
@@
 In addition to fragment matching, we will also discuss approaches that solve the correlated problems of fragment segmentation and fragment matching together.
-Similarly to single-piece analysis, we will cover both non-deep learning approaches and deep learning approaches.
+Similarly to single-piece analysis, we will cover both procedural approaches and deep learning approaches.
 Template shape space (Sec.
@@
 7).
-After discussing technical approaches towards solving the reassembly problem, we will survey various applications of the reassembly problem.
-We will discuss five areas of fractured object reassembly application, including bone reduction, restoring fragments of thin-shells, walls, and building, and reassembly problems in science.
-Finally, we discuss the relevant problem of assembling parts to form complete objects and assembling objects to form 3D scenes.
-Datasets and software packages (Sec.
-8).
-We discuss publicly available datasets and open source software packages.
-Datasets include early small-scale datasets and recent large-scale real- and synthetic datasets.
+After discussing technical approaches to solving the reassembly problem, we will discuss various applications of the reassembly problem.
+We will discuss five areas of fractured object reassembly application, including bone reduction, restoring thin-shell, planar, and stone fragments, and reassembly problems in science.
+Finally, we discuss the relevant problem of assembling parts to form complete objects.
+Datasets and open source (Sec. 8). We discuss publicly available datasets and open source packages. Datasets include early small-scale datasets and recent large-scale real- and synthetic datasets.
 Future directions (Sec.
 9).
-The last paper of this survey summarizes existing approaches and discusses future directions.
+The last section of this survey summarizes existing approaches and discusses future directions.
 We will emphasize how to advance existing reassembly problems by developing better learning approaches, focusing on challenges in representations, data, and utilizing foundation models.
@@
 3.1 Reconstruction
-Physical fragments must be captured and reconstructed into digital models before reassembling.
-This section discusses how fragmented pieces can be captured using various sensors, each requiring specific reconstruction techniques.
-Laser scanning.
-As shown in Figure 2 (Top-Left), laser scanners capture the 3D geometry of fragments by emitting laser beams and measuring the time or phase difference of the reflected light.
-They allow for high-resolution geometry capture, achieving a physical distance of approximately $0.01$ mm between each captured point.
-The acquisition efficiency for capturing a piece, depending on the resolution selected, laser emission frequency, and the number of emitted laser lines, typically ranges from minutes to hours.
-For example, scanning a $10$ cm object using the low-frequency NextEngine scanner ($50K$ points per second) would take about an hour [18].
-However, with high-frequency HandySCAN ($1.8M$ points per second), the process would only take a few minutes [19, 20].
-Laser scanning is usually desirable for capturing high-quality geometry, particularly in outdoor environments.
-However, since it does not capture color information, other capturing modalities should be combined if chromatic data is needed.
-A significant drawback of these scanners is their high cost, though rental options are available.
-There are several representative datasets captured by laser scanning.
-Stanford University used laser scanners to collect the “Forma Urbis Romae” dataset as part of the Digital Michelangelo Project, capturing high-resolution 3D shapes of marble fragments for digital preservation and research [21].
-Another significant dataset is generated through the “Reconstructing Thera Frescoes” project, where low-power laser scanning is employed to ensure accurate data capture while also protecting fragile frescoes from potential damage during the scanning process.[22]
-Photogrammetry capturing.
-Color images captured by regular cameras can also be used to capture fragment geometry and texture; and recent studies have found that geometry reconstruction quality can be comparable to laser scanning, with a resolution of $0.15$  mm and an accuracy of $0.13$  mm [23, 24].
-State-of-the-art multi-view stereo (MVS) algorithms reach pixel level accuracy, namely, $90\%$ reconstructed points have accuracy within about one pixel [25].
-Therefore, the geometry quality can be further improved by capturing detailed images from a short range.
-As shown in Figure 2 (Top-Right), photogrammetry is well suited for field data collection because it does not require specialized skills to operate a scanner, is affordable and easy to set up.
-Designed for field collection without electricity, [26] proposed a rig to do photogrammetrical documentation of artifacts.
-Data collection for this protocol averaged approximately $20$ minutes, with a series of overlapping photographs (approximately $100$ per artifact) to be taken of the artifact from all angles [27].
-To obtain even higher throughput, the batch method has been proposed by [24], achieving a throughput of about $730$ fragments per day.
-However, photogrammetry is sensitive to light and is not suitable for transparent material.
-The lighting needs to be closely controlled using two overhead fluorescent lamps or LEDs, which minimize the amount of shadows and regularized the amount of light across the surface[18].
-Note that a list of “non-friendly” materials for photogrammetric purposes, such as water, glass, varnish, steel, and many more, can be found in [28].
-In addition, selecting an appropriate focal length is essential to achieve optimal results in the digitalization process [29].
-Wang et al.
-proposed a comprehensive dataset consisting of batches of 3D sherd scans, typically pottery fragments, accompanied by auxiliary information that aids in the alignment and reconstruction of the original artifacts through model registration techniques [24].
+Converting physical fragments into digital 3D models involves two key stages: digitization through various scanning technologies and registration to create complete 3D models.
+Although scanning choices influence data quality, the registration process shares fundamental algorithmic similarities with the assembly task itself.
+This section discusses how fragmented pieces can be captured using sensors, each of which requires specific reconstruction techniques.
+Scanning Technologies. Several technologies are available for fragment digitization (Figure 3):
+Laser Scanning captures high-resolution geometry ($\sim 0.01$ mm) by measuring reflected laser beams.
+High-frequency systems like HandySCAN ($1.8M$ points/second) enable rapid capture [19, 20], while lower-frequency options like NextEngine provide more affordable alternatives.
+Laser scanning excels in outdoor environments and has enabled major digitization efforts like Stanford’s “Forma Urbis Romae” dataset [21] and the “Reconstructing Thera Frescoes” project [22].
+However, the lack of color capture and high equipment costs are notable limitations.
+Photogrammetry achieves resolution comparable to that of laser scanning ($\sim 0.15$ mm) [23, 24] using a regular camera.
+State-of-the-art multi-view stereo algorithms achieve pixel-level accuracy for 90% of reconstructed points [25].
+This approach is particularly suited for field work due to its low cost and minimal equipment requirements.
+Porter et al.
+[26] designed a specialized rig for field collection without electricity, requiring approximately 20 minutes and 100 photographs per artifact [27].
+Although sensitive to lighting conditions and require controlled illumination [28], recent protocols have demonstrated a high throughput of 730 fragments per day [24].
+However, photogrammetry struggles with certain materials such as water, glass, and steel [29], and requires careful selection of focal length [30].
+Structured Light balances speed and accuracy ($\sim 0.1$-$0.2$ mm) by projecting patterns onto objects [31].
+It can process 60 pottery fragments per hour [32] or 5-6 reflective artifacts per hour [33].
+While capturing both geometry and color, it works best indoors and may struggle with shiny or dark surfaces.
+Notable applications include the work by Fan et al.
+[34].
+CT Scanning is uniquely capable of capturing internal structures [35, 36] at micron-scale resolution ($1$-$500$ microns) [37].
+It has been applied successfully to rocks [38], stones [39], fossils [40], bones [41], and teeth [42].
+Recent work demonstrates batch processing of 220 artifacts in two hours at 140-micron resolution [39].
+Important datasets include skull scans [43] and detailed anatomical structures [44].
 Figure 3: Four types of sensors for digitizing fragment shapes.
@@
 (Bottom-right) CT Scanning [46].
-Structured light capturing.
-As shown in Figure 2 (Bottom-Left), The main advantages of this technology are speed, making it suitable for applications where efficiency is crucial [32].
-This method operates by projecting structured light onto an object and then analyzing the reflected light to determine the shape and dimensions of the object.
-The resolution of structured light scanners, which is around $0.1\sim 0.2$ mm, is slightly inferior to that of laser scanners.
-However, it can generally obtain better geometric detail and resolution in small details than digital photogrammetry.
-The operation time can range from 1 to 10 minutes, depending on the optical properties of the scanned object.
-In [33], 60 pottery fragments can be scanned per hour at the desired resolution.
-For some reflective pieces, such as flint, an experienced operator produced complete digital 3D images at the rate of $5\sim 6$ artifacts per hour [34].
-Compared with laser, structured light can capture color information, but may struggle with shiny or dark surfaces.
-Therefore, they are typically well-suited for scanning small parts that require color and texture details, and for use in indoor environments.
-A representative project in this category is [35].
-CT scanning.
-CT scanners are extremely useful in piece reconstruction because they allow 3D scanning not only of the outer surfaces of an object but also hidden internal structures [36, 37].
-Unlike outer surface geometry, internal structures of the pieces are difficult to capture using the aforementioned scanning technology that relies on light reflection.
-The resolution is typically approximately $1\sim 500$ microns, a level of precision capable of capturing the fine topology [38].
-To date, numerous studies have used CT to image rocks [39], stones [40], fossils [41], bone [42] and teeth [43].
-Even though a high-resolution CT-scan can take several hours, the duration needed to scan a single part is manageable.
-[40] were able to scan up to 220 lithics at once in a relatively short period of time (ca.
-2 hours) at a resolution of $140$ microns.
-Their tests suggested that this resolution is optimal for scanning many artifacts simultaneously while maintaining accuracy in the 3D shapes.
-The skull dataset [44] comprises high-resolution CT scans of both damaged and normal skulls, aiding in the study and reassembly of cranial anatomy.
-Another dataset [45] consists of high-resolution CT scans generated through helical scanning, providing detailed 3D reconstructions of anatomical structures stored in DICOM format.
 Registration and 3D reconstruction.
@@
 When reference markers are possible, such as small disc-shaped stickers [19] or unevenly spaced grid lines drawn on a turntable [28], they facilitate point cloud registration by serving as common reference points.
-When using such markers is infeasible, markerless registration such as Iterative Closest Point (ICP) algorithm and its variants are utilized to align the point clouds [33].
-The quality of the initial setup in the Iterative Closest Point (ICP) algorithm is crucial, as it significantly influences the convergence rate and precision of the algorithm [46].
-Therefore, careful manual annotations are essential for the optimum outcomes.
+When using such markers is infeasible, markerless registration such as the Iterative Closest Point (ICP) algorithm and its variants are utilized to align the point clouds [32].
+A good initial alignment is essential for the ICP algorithm, as it directly affects how quickly and accurately the algorithm can find the final alignment [47].
+Therefore, careful manual annotations are essential for optimum outcomes.
 In order to improve registration efficiency, [24] devised an automated and robust registration technique that leverages the bilateral boundary ICP, offering reliable initial relative pose estimates to initiate the ICP.
-After forming the dense point cloud, a mesh modeling algorithm, such as marching cubes [47] or Poisson reconstruction [48] is used to generate the surface, In contrast to the methods that record the scans before reconstructing the surface, [49] integrates both processes simultaneously, resulting in a noise-free and watertight solid model.
+After forming the dense point cloud, a mesh modeling algorithm, such as marching cubes [48] or Poisson reconstruction [49] is used to generate the surface.
+Unlike traditional methods that separate scanning and surface reconstruction, [50] combines these steps into a single process to create cleaner and more complete 3D models.
 Using joint registration, the method significantly improves reconstruction accuracy, ensuring better consistency when combining different views.
 The experimental results indicate that it can perfectly reconstruct the surfaces of archaeological fragments.
-This approach was used to reconstruct the fragments of the Octogan project [50].
+This approach was used to reconstruct the fragments of the Octagon project [51].
 Batch processing in scanning improves efficiency and reduces the need for manual intervention.
 This approach is widely adopted in large-scale applications.
-[35] proposed a system to automatically scan multiple 3D objects using view and path planning algorithms, minimizing manual intervention.
+[34] developed an automated system that can scan multiple 3D objects by intelligently planning camera views and scanning paths, reducing the need for human input.
 It enhances scanning efficiency and quality with structured light scanning and calibrated positioning, making it ideal for large-scale applications like archaeological documentation.
@@
 The fast geometry-based matching further optimizes efficiency, allowing the system to process approximately 10 fragments per hour.
-[24] presented another batch system to scan two sides of fragments and developed tailor registration and reconstruction techniques to process the paired scans.
+[24] presented another batch system to scan two sides of fragments and developed custom registration and reconstruction techniques to process the paired scans.
 3.2 Segmentation
-Surface segmentation is a customary preprocessing step in the analysis of individual fragments for reassembly tasks, aimed at breaking down complex surface geometries into simpler, meaningful regions.
+Surface segmentation is a common preprocessing step in the analysis of individual fragments for reassembly tasks, aimed at breaking down complex surface geometries into simpler, meaningful regions.
 Typically, this involves distinguishing between the original and fracture surfaces, facilitating easier analysis and subsequent matching with other fragments to achieve successful reassembly.
-Geometry-based method.
+Geometry-based methods.
 Geometry-based surface segmentation methods rely on the intrinsic geometric properties of the fragment’s surface, such as curvature, edges, and other spatial features.
@@
 Several geometry-based methods have been developed for surface segmentation.
-For instance, [51] employ a simple region-growing algorithm primarily based on normal vectors, along with a facet segmentation threshold, to initially segment the original object into distinct surface regions.
+For example, [52] employs a simple region-growing algorithm based primarily on normal vectors, along with a facet segmentation threshold, to initially segment the original object into distinct surface regions.
 Following this segmentation, bumpiness is used to detect fractured surfaces.
@@
 In addition to these, various other clustering and graph segmentation algorithms, such as K-means clustering [56], hierarchical clustering [57], mean-shift clustering [58], spectral clustering [59] and random walks [60], can also be adapted for fractured surface segmentation.
-These techniques are well-established, relying on explicit rules and geometric properties.
-They tend to be more interpretable and less data-intensive but are limited by surface complexity and noise.
-Figure 3: (Top) Non-deep learning methods for fracture surface segmentation [51]. (Bottom) Deep learning methods for fracture surface segmentation [60].
+These techniques leverage explicit geometric rules and properties, making them interpretable and efficient with limited data.
+However, their performance may decrease when handling complex surfaces or noisy input.
+Figure 4: (Top) Procedural methods for fracture surface segmentation [52]. (Bottom) Deep learning methods for fracture surface segmentation [61].
 Figure 5: A representative multi-piece shape analysis framework presented in [53].
-Pair-wise matching use use patch features extracted from fracture surface and segment features extracted from fracture surface boundaries.
+Pair-wise matching use patch features extracted from the fracture surface and segment features extracted from fracture surface boundaries.
 A voting-based approach is used to find consistent correspondences between geometric features.
-Matching among pieces employ a greedy approach.
-Learning-based method.
-Learning-based surface segmentation techniques have gained traction in recent years due to their ability to automatically learn patterns and features from data, without the need for explicit geometric rules.
-These methods, particularly those involving deep learning, are transforming how surface segmentation is performed in fragment reassembly tasks.
+Matching among pieces employs a greedy approach.
+Learning-based methods.
+Learning-based surface segmentation techniques have emerged as a promising direction, as they can automatically discover patterns and features from data without requiring hand-crafted geometric rules.
+These methods offer new possibilities for surface segmentation in fragment reassembly tasks.
 Although some learning-based fractured object reassembly methods take the entire fragments as input for analysis, bypassing prior surface segmentation, there are still methods that decouple the reassembly pipeline, retaining segmentation to achieve better matching and reduce computational costs.
@@
 The initial segmentation process can be conceptualized as a binary classification task.
-For example, [60] utilize a convolutional neural network (CNN) to learn the labels of ridge and non-ridge cells (see Figure 3(Bottom)), facilitating the extraction of the fracture margins from the fragments.
-Subsequently, fractured surface can be well segmented by trimming the fragment models along these margins.
-[61] employ Multi-Layer Perceptron (MLP) layers combined with a sigmoid activation function to predict the confidence in identifying a fracture point.
+For example, [61] utilize a convolutional neural network (CNN) to learn the labels of the ridge and non-ridge cells (see Figure 4(Bottom)), facilitating the extraction of the fracture margins of the fragments.
+Subsequently, fractured surfaces can be well segmented by trimming fragment models along these margins.
+[62] employ Multi-Layer Perceptron (MLP) layers combined with a sigmoid activation function to predict confidence in the identification of a fracture point.
 If we generalize segmenting fractured surfaces to 3D semantic segmentation, we can also leverage advanced general-purpose graph learning algorithms like PointNet [63], PointNet++ [64] and DGCNN [65] to perform segmentation.
 These powerful frameworks offer robust solutions for intricate 3D data analysis, and they will also be mentioned in the next section.
-These approaches, particularly deep learning, offer more flexibility and can handle complex surfaces and noisy data better.
-However, they require large datasets and significant computational power.
+These frameworks provide effective tools for analyzing complex 3D data, as we will discuss further in the next section.
+While these approaches can potentially handle challenging cases with complex surfaces and noise, they typically require large training datasets and substantial computational resources.
 4 Multi-Piece Analysis
@@
 4.1 Pairwise Matching
-Rigid scan maching.
+Rigid scan matching.
 A relevant problem of pairwise piece matching is rigid scan matching, which finds the relative rigid pose between two depth scans.
-Rigid scan matching is an important step in 3D geometry reconstruction from depth scans.
+Given two point sets $P$ and $Q$, rigid scan matching aims to find a rotation $R\in SO(3)$ and translation $t\in\mathbb{R}^{3}$ that minimize:
+$\min_{R,t}\sum_{i=1}^{n}\min_{j}||Rp_{i}+t-q_{j}||^{2}$ (1)
+where $p_{i}\in P$ and $q_{j}\in Q$ are points from the respective scans.
 A standard pipeline of rigid scan matching consists of feature extraction, feature matching, and rigid pose fitting.
 Early approaches use hand-crafted geometric descriptors, e.g., SpinImage [66] and Integral invariant [67] features.
-These features usually only encode local information; therefore, the resulting feature matches have a lot of false positives.
+These features usually only encode local information; therefore, the resulting feature matches have many false positives.
 The rigid pose fitting procedure addresses this issue by enforcing the rigidity constraint, which finds a subset of feature matches that share a rigid transformation.
 Popular rigid pose fitting approaches include RANSAC [68], general Hough transform [69, 70, 71], and their variants [72, 73].
-Another category of methods formulates rigid scan matching as spectral matching, which computes a consistency matrix among candidate feature matches.
+Another category of methods formulates rigid scan matching as spectral matching [74], which computes a consistency matrix among candidate feature matches.
 The elements of this consistency matrix encode whether distances and angles are preserved between the corresponding pairs of correspondences.
-Rigid pose fitting is formulated as computing a sub-matrix in which all elements are consistent, and this is done using spectral techniques.
+Rigid pose fitting is formulated as computing a sub-matrix in which all elements are consistent, and is done using spectral techniques.
 More recent approaches develop deep learning techniques for rigid scan matching.
-These approaches employ deep neural networks to extract features and develop differentiable layers for fitting rigid poses.
+These approaches employ deep neural networks to extract features and develop differentiable layers to fit rigid poses.
 Examples include differentiable RANSAC and attention-based rigid regression.
 Representative methods in this category include DCP [75], PNetLK [76], FCGF [77], Predator [78], RoReg [79], and GeoTF [80].
-These approaches show significantly improved performance on benchmark datasets where the domain gap between the testing set and the training set is small.
-However, it is still an open question to develop generalizable deep learning approaches for rigid scan matching.
-Non deep fragment matching.
+These approaches demonstrate promising results on benchmark datasets with similar training and testing conditions.
+Developing approaches that generalize well to diverse real-world scenarios remains an important research direction.
+Procedural fragment matching.
 Similar approaches have been developed for rigid fragment matching, which also combines feature extraction, feature matching, and rigid pose fitting.
 However, rigid fragment matching and rigid scan matching have several notable differences.
-These differences come from several aspects, including where to extract features and what constraints are induced from feature matching.
+These differences come from several aspects, including where to extract features and what constraints are induced by feature matching.
 Another important difference is the robustness of the extracted features, the feature descriptors, and the matching procedure, as the surfaces of the fragments may be severely damaged or there are tiny fragments between two fragments, causing geometric differences.
@@
 Robustness has been a major concern when extracting features from fracture surfaces.
-Therefore, many features introduced in the literature are patch features that are introduced to improve robustness.
+Therefore, many features introduced in the literature are patch features that are developped to improve robustness.
 [53] use patches defined by sub-level sets of the integral invariant descriptors and summarize geometric descriptors derived from performing principal component analysis of patch shapes.
@@
 A key issue in matching fracture surfaces is developing descriptors that can be used to identify geometric patches that can be tucked together [86].
-A recent paper [61] also addresses this issue by introducing a primal-dual descriptor in deep learning approach.
-Many papers focus on extract from boundary curves of fracture surfaces [84, 81].
+A recent paper [62] also addresses this issue by introducing a primal-dual descriptor in the deep learning approach.
+Many papers focus on extracting boundary curves of fracture surfaces [86, 83].
 Such features are more robust in the setting where the areas of the fracture surfaces are small, making it difficult to compute and extract geometric descriptors.
+[87] performs (potentially partial) contour matching of the break boundaries to constrain the pairwise surface registration.
 [88] analyze and simulate fracture patterns of Theran wall paintings.
@@
 [92] introduced a fast algorithm for 2D fragment assembly based on a partial EMD distance.
-[90] introduced a descriptor based on the convex/concave information of a point on the boudnary curves of the fractured surface.
+[93] introduced a descriptor based on the convex/concave information of a point on the boundary curves of the fractured surface.
 The similarity between two fractured surfaces is calculated based on the spin images of the point of the feature curve and the distance and normal deviation between the two feature curves to find the matching fractured surface.
@@
 Second, the pairwise reassembly probability is calculated using LCD and SDD descriptors; then, collision detection is performed to eliminate the incorrect overlapping pairs.
-Several approaches rely on features on the original surfaces for matching (see Figure 5).
+Several approaches rely on the features on the original surfaces for matching (see Figure 6).
 Early work by [97] discussed how to use original surface features for geometric matching.
@@
 They proposed confidence scores on potential matches.
-Several approaches define scoring functions over the space of transformations and develop global search strategies to global optimal solutions.
+Several approaches define scoring functions over the space of transformations and develop global search strategies for global optimal solutions.
+[100] uses simulated annealing to search the space of the source to transformation, while employing hardware-accelerated rasterization to extract surface distances based on depth maps and compute a similarity score using distance gradients.
 [101] introduced an approach that focuses on wall-painting fragments and applies a discrete search of relative poses to identify global solutions to reassembly.
@@
 In addition to geometric matching, another difference between scan matching and fragment matching is that the matches should be free of penetrations.
-This constraint is easy to integrate for RANSAC-based approaches [67, 98].
+This constraint is easy to integrate for RANSAC and stochastic global optimization approaches [68, 102].
 An open question is how to enforce this under global optimization approaches such as spectral matching.
@@
 Early work includes Deepzzle [107], which studied image reassembly, a special case of fractured object reassembly.
-Deepzzle studied how to handle noisy inputs, e.g., erosion, missing pieces, and fragments of other objects.
-Thanks to the emergence of 3D benchmark datasets for fractured object reassembly, there are increasing interests in developing deep learning approaches to solve the 3D fractured object reassembly.
-[104] proposed MatchMakerNet (see Figure 9(Left)), a network architecture designed to automate the pairing of object fragments for reassembly.
+Deepzzle explored methods to handle real-world challenges such as erosion, missing pieces, and the presence of unrelated fragments.
+Due to the emergence of 3D benchmark datasets for fractured object reassembly, there is increasing interest in developing deep learning approaches to solve the 3D fractured object reassembly.
+[108] proposed MatchMakerNet (see Figure 10(Left)), a network architecture designed to automate the pairing of object fragments for reassembling.
 The approach takes two point clouds as input and its network leverages graph convolution alongside a simplified version of DGCNN for prediction.
-However, most existing approaches have focused on feature learning and incorporating rigidity constraints and problem-specific knowledge deserves future research.
+Current approaches have primarily focused on feature learning while incorporating rigidity constraints and domain-specific knowledge remains an important direction for future research.
+Bridging deep learning methods with practical applications continues to be an active area of investigation.
 Figure 8: Enforcing non penetration via constrained optimization [53, 103, 104, 105].
@@
 Optimization-based based approaches.
-The difference between multiple-piece matching and pairwise matching is that there is a cycle-consistency constraint [105, 106, 107] among cycles of fragments.
-Specifically, when composing relative poses along a cycle of fragments, we should arrive at the identity map.
-If not, then at least one of the relative poses is incorrect.
-This cycle-consistency constraint can be used to prune incorrect pair-wise matches computed between pairs of objects in isolation.
+The difference between multiple-piece matching and pairwise matching is the cycle-consistency constraint [109, 110, 111] among cycles of fragments.
+Specifically, when composing relative transformations $T_{ij}$ along a cycle of fragments $(i_{1},i_{2},...,i_{k})$, we should have:
+$T_{i_{1}i_{2}}T_{i_{2}i_{3}}...T_{i_{k-1}i_{k}}T_{i_{k}i_{1}}=I$ (2)
+where $I$ is the identity transformation.
+If this constraint is violated, then at least one of the relative poses along the cycle must be incorrect.
+This cycle-consistency constraint is used to prune incorrect pair-wise matches computed between pairs of objects in isolation.
 In the literature on scan matching, early approaches enforce combinatorial optimization to search for consistent matches.
@@
 On the other hand, combinatorial optimization approaches are typically greedy and can easily return sub-optimal solutions.
-Another category of approaches formulate multi-piece matching as Markov-random field (MRF) inference.
+Another category of approaches formulates multi-piece matching as Markov-random field (MRF) inference.
 This is formulated as sampling a discrete set of candidate poses of each object and model pairwise matches as pairwise potentials to optimize the best candidate of each object.
@@
 A relevant approach is [116], which presents a linear programming (LP) relaxation to the MRF formulation on the image-based square patch puzzle problem.
-This MRF formulation was later applied for in MRF-SFM [113, 114] to solve the structure-from-motion problem.
-The nice properties of this MRF formulation is that it possesses better solvers that offer improved solutions compared to greedy approaches.
-Moreover, it is easy to incorporate the non-penetration constraint, i.e., if the a pair of fragments penetration under two candidate poses, then the corresponding pairwise potential is infinite.
+This MRF formulation was later applied in MRF-SFM [117, 118] to solve the structure-from-motion problem.
+The nice properties of this MRF formulation are that it possesses better solvers that offer better solutions compared to greedy approaches.
+Moreover, it is easy to incorporate the non-penetration constraint, i.e., if two fragment penetrates under two candidate poses, then the corresponding pairwise potential is infinite.
 The limitation of this MRF formulation is that it requires many pose samples to obtain accurate solutions.
-Modern multi-piece matching typically formulates low-rank matrix recovery by establishing the equivalence between the cycle-consistency constraint and the fact that the matrix that encodes pairwise maps in block is low-rank, c.f.
-[115].
+Modern multi-piece matching typically formulates low-rank matrix recovery by establishing the equivalence between the cycle-consistency constraint and the fact that the matrix that encodes pairwise maps in the block is low-rank, c.f., [119].
 The optimization approaches fall into categories of semidefinite programming relaxation [120, 121, 119, 122, 123, 124, 125], spectral techniques [126, 127, 111, 128, 129, 130, 131, 132, 133, 134, 135, 136], and non-convex optimization [137, 138, 139, 140, 141, 142, 143].
@@
 Figure 10: Recent advances on deep fragment matching. (Left) Techniques that extend rigid matching [108]. (Middle) Transformer-based techniques [145]. (Right) Equivariant techniques [146].
-Applications of these multi-object matching approaches in the context multi-fragment reassembly have been limited.
+Applications of these multi-object matching approaches in the context of multi-fragment reassembly have been limited.
 Most approaches follow greedy approaches [53, 147, 145].
 There are a few exceptions.
-[144] introduced an unsupervised genetic algorithm to reassemble wall painting fragments.
+[148] introduced an unsupervised genetic algorithm to reassemble fragments of wall paintings.
 The approach evolves a pool of partial reconstructions that grow through recombination and selection over the course of generations.
 The authors introduced a novel algorithm for combining partial reconstructions that is robust to noise and outliers, and a new selection procedure that balances fitness and diversity in the population.
-[86] detect loop in the pose graph and performs loop closure to rectify pose errors.
+[89] detect loop in the pose graph and perform loop closure to rectify pose errors.
 [144] enforce global Consistency in the Automatic assembly of fragmented artifacts.
@@
 [96] introduced a global optimal reassembly solution is obtained by iterative graph optimization with the constrains of the loop-closures and overlap restrictions.
-One critical difference between multi-scan matching and multi-fragment matching is that the underlying graph of the fragments is much more sparse than the graph in multi-scan matching, as scans have significantly overlap, while fragments do not overlap.
+One critical difference between multi-scan matching and multi-fragment matching is that the underlying graph of the fragments is much more sparse than the graph in multi-scan matching, as scans have significant overlap, while fragments do not overlap.
 Deep pose regression based approaches.
@@
 [145] introduced a recurrent approach that aligns and merges fragments into larger groups.
-The key contribution is a diffusion model simultaneously denoises the 6-DoF alignment parameters of the fragments, and a transformer model verifies and merges pairwise alignments into larger ones, whose process is repeated iteratively.
+The key contribution is a diffusion model that simultaneously denoises the 6-DoF alignment parameters of the fragments, and a transformer model that verifies and merges pairwise alignments into larger ones, whose process is repeated iteratively.
 [153] introduced an approach that identifies critical stones among the input fragments for efficient computation of the reassembly procedure.
@@
 [154] use a hybrid attention module to model and reason complex structural relationships between fragment patches.
-The model has intra- and inter-attention layers, enabling capturing of crucial contextual information within fragments and relative structural knowledge across fragments.
+The model has intra- and inter-attention layers, enabling the capturing of crucial contextual information within fragments and relative structural knowledge across fragments.
 The model is enhanced with an adjacency aware hierarchical pose estimator.
-People have developed equivalent networks [151] for point cloud learning to solve fractured object reassembly.
+Researchers have developed equivalent networks [155] for point cloud learning to solve fractured object reassembly.
 [146] developed equivariant and invariant encoders of each part to compute equivariant and invariant correlation modules to aggregate extracted signals across all parts, which are used to regress the pose of each input part.
@@
 BITR can not only handle non-overlapped PCs, but also guarantee robustness against initial positions.
-Extension of SE(3) equivariance.
-[153] introduced DiffAssemble, which follows Diffusion Probabilistic models formulations.
+[157] introduced DiffAssemble, which follows Diffusion Probabilistic model formulations.
 DiffAssemble models a Markov chain in which it injects noise into the position and orientation of the pieces.
@@
 The denoising network is an attention-based GNN that takes as input a graph where each node contains an equivariant feature that describes a particular piece and its position and orientation.
-Similarly, [154] and [141] treated the reassembly task as a denoising task and introduced FragmentDiff and PuzzleFusion++, a method that applies diffusion denoising through transformer to predict the pose parameters of each fragment based on their global feature correlations and learned pose priors.
+Similarly, [158] and [145] treated the reassembly task as a denoising task and introduced FragmentDiff and PuzzleFusion++, methods that apply diffusion denoising through the Transformers to predict the pose parameters of each fragment based on their global feature correlations and learned pose priors.
+While these deep learning approaches show promising results, bridging the gap between synthetic training data and real-world applications remains an important challenge.
+Real fragments often exhibit erosion, weathering, or missing pieces - conditions that are difficult to fully capture in synthetic training data.
+Understanding how to better generalize to these practical scenarios is an active area of research.
 4.3 Joint Segmentation and Matching
-Figure 10: Joint matching and segmentation [61].
-Figure 11: (a) Early work on linear shape spaces [155]. (b) Deep deformable shape generators [156]. (c) Implicit shape generators [157, 158, 159]. (d) Point cloud shape generators [160].
+Figure 11: Joint segmentation and matching [62].
 As discussed above, an important problem in fragment analysis is segmenting the fracture surface.
@@
 [161, 162, 122, 163] introduced the functional map framework which allows us to optimize consistent segmentations by optimizing a function sub-space of each shape.
-In the context of fractured object reassembly, the only approach that explicitly models this objective is JIGSAW [61] (see Figure 10), It extracts point-wise features using transformers, which are used to predict point labels of original or fracture surfaces and perform point-wise matching.
-Joint segmentation and matching are achieved using shared deep point-wise feature descriptors.
+In the context of fractured object reassembly, JIGSAW [62] (see Figure 11) pioneered a deep learning approach for joint segmentation and matching.
+It effectively utilizes transformer architecture to extract point-wise features and proposed primal-dual descriptors for complementary matching.
+JIGSAW also demonstrates how learning approaches can be effectively combined with classical geometric algorithms.
+By integrating learned point-wise features with RANSAC and transformation synchronization methods, it achieves robust pose estimation.
+This hybrid approach leverages both the pattern recognition capabilities of deep learning and the geometric guarantees of classical methods, suggesting that combining complementary techniques may lead to more effective solutions.
 5 Template Shape Space
+Figure 12: (a) Early work on linear shape spaces [164]. (b) Deep deformable shape generators [165]. (c) Implicit shape generators [166, 167, 168]. (d) Point cloud shape generators [169].
 In many cases, we do not have prior knowledge of the exact shape of the underlying complete object.
@@
 Early approaches (see Figure 12(a)) include Morphable faces [170] and SCAPE [69], which focus on 3D shape spaces with the same topology.
-Recent examples along this line include SMPL [155] for humans, SMAL [167] for animals, MANO [168] of hands, and 3DMM [169] for faces, which are learned from larger datasets and therefore exhibit significantly improved performance in solving inverse problems.
-These models decouple shape variations and pose variations.
-Shape variations are derived from PCA analysis.
-Pose variations, e.g., in SMAL and SMPL are encoded using bones.
-A limitation of these approaches is that require consist meshes to learn the shape prior.
+Recent examples along this line include SMPL [164] for humans, SMAL [171] for animals, MANO [172] of hands, and 3DMM [173] for faces.
+These models, trained on larger datasets, demonstrate significantly improved performance in solving inverse problems.
+They decompose shape variations from pose variations - shape variations are captured through PCA analysis, while pose variations (as in SMAL and SMPL) are encoded using skeletal structures.
+However, a key limitation of these approaches is that they require training data with consistent mesh topology, meaning that all meshes must share the same vertex structure.
 Recent approaches have focused on using deep neural networks to learn generative models, which are superior to linear models.
@@
 Despite significantly improved performance from linear models, these approaches still require consistent inter-shape correspondences as input for learning.
-DeepSDF [157] and follow-up work show that shape generative models can be learned from raw data points with correspondences.
+DeepSDF [166] and follow-up work show that shape generative models can be learned from raw data points without correspondences.
 SALD [179] shows that MLP-based implicit surfaces can also encode deformable shapes, but generalization of shapes that undergo large pose changes is very limited.
@@
 However, one limitation of using the implicit representation is that it is very difficult to compute the correspondences between the learned template and the input fragments.
-Figure 12: Using prior knowledge from a complete object model to aid reassembly.
-(a) Local-global alignment with a selected explicit template [189], such as a skull.
-(b) Implicit learned prior knowledge from a large-scale 3D object dataset [190].
-(c) A global evaluation using neural network as complete object prior [191, 192].
-Table 2: Category of the different complete object priors with their descriptions and limitations.
-Categories Descriptions Limitations
-Local-global Alignment Shape prior provides reasonable initial reassembly Applicable to limited types of reassembly objects
-Global Evaluation Regularize the quality of the reassembled result Hard to find a proper function for each reassembly
-One way to address this issue is to learn generative under explicit representations such as point clouds and meshes.
+One way to address this issue is to learn generative models under explicit representations such as point clouds and meshes.
 Many approaches have shown promising results for point cloud generators for man-made shapes [193, 194, 195, 169, 196].
@@
 6.1 Complete object priors
-In addition to feature-based reassembly, a complete object model from the same or a similar class can provide valuable prior knowledge to facilitate reassembly.
-For example, when reassembling objects such as skulls, pottery, and architectural elements, complete models are often available and can serve as templates to infer the initial arrangement of fragments.
-In recent data-driven techniques, these global geometric priors can also be learned from collected data sets using neural networks.
+Beyond purely feature-based approaches, using complete objects as templates can significantly boost the reassembly process.
+This is particularly valuable when working with well-documented objects like skulls, pottery, or architectural elements, where intact reference models often exist.
+Recent advances in machine learning have also made it possible to learn these geometric priors from collections of similar objects, offering more flexible template-based solutions.
+Figure 13: Using prior knowledge from a complete object model to aid reassembly.
+(a) Local-global alignment with a selected explicit template [200], such as a skull.
+(b) Implicit learned prior knowledge from a large-scale 3D object dataset [201].
+(c) A global evaluation using neural network as complete object prior [202, 203].
+Table 2: Category of the different complete object priors with their descriptions and limitations.
+Categories Descriptions Limitations
+Local-global Alignment Shape prior provides reasonable initial reassembly Applicable to limited types of reassembly objects
+Global Evaluation Regularize the quality of the reassembled result Hard to find a proper function for each reassembly
 This section discusses how a complete object model or global prior can be used to facilitate object reassembly.
@@
 In contrast, as discussed in Sec.
-6.3, recent learning-based methods tend to train neural networks to build implicit shape priors from datasets, and then train/optimize the encoding of both local fragments and global template so that their latent codes can infer the matching and alignment of the fragments.
-In classical geometric methods, to support effective local-global alignment, the explicit shape template is often selected as an average or typical shape, or a best-matched object retrieved from a database.
-Using a typical object as prior shape is common in skull and cultural heritage reassembly tasks.
+6.3, recent learning-based methods tend to train neural networks to build implicit shape priors from datasets and then train/optimize the encoding of both local fragments and global template so that their latent codes can infer the matching and alignment of the fragments.
+In classical geometric methods, to support effective local-global alignment, the explicit shape template is often selected as an average or typical shape or a best-matched object retrieved from a database.
+Using a typical object as the prior shape is common in skull and cultural heritage reassembly tasks.
 For example, [204, 205, 200] used a standard human skull as $P$ to provide alignment guidance.
@@
 In the archaeological and architectural reconstruction of a ruined monument task, to construct the most similar geometry as prior $P$, [51, 210] conducted a careful study of historical texts that mention the building under consideration.
-Given the selected shape prior $P$, as for the second factor, feature extraction computes the keypoints and their descriptors on both fragments and the global prior, such as depth and axial symmetry descriptor [203], salient points [204], parabolic contour [87], spin images [200, 201], Signature of Histograms of OrienTations (SHOT) descriptor [189].
+Given the selected shape prior $P$, as for the second factor, feature extraction computes keypoints and their descriptors on both fragments and the global prior, such as depth and axial symmetry descriptor [207], salient points [208], parabolic contour [90], spin images [204, 205], Signature of Histograms of OrienTations (SHOT) descriptor [200].
 With the extracted keypoints and features, fragment transformations are computed to roughly assemble the input parts to the template.
-Researchers developed various alignment methods using keypoint correspondences, such as iterative closed points (ICP) [200], geometric distance [201], local affine-invariant 3D moments [87], Earth mover’s distance [204], and RANSAC algorithm [189].
+The researchers developed various alignment methods using keypoint correspondences, such as iterative closed points (ICP) [204], geometric distance [205], local affine-invariant 3D moments [90], Earth mover distance [208], and the RANSAC algorithm [200].
 Implicit prior for local-global alignment is learned from an annotated 3D dataset to solve the partial matching from the fragments to the global shape.
-The shape prior is encoded in a neural network, which is trained on a 3D segmentation dataset, that the weights are adjusted to learn the semantic relationship between the fragments and the complete objects.
+The shape prior is encoded in a neural network, which is trained on a 3D segmentation dataset, and the weights are adjusted to learn the semantic relationship between the fragments and the complete objects.
 In the inference stage, given a set of parts $X$, trained networks are used as encoders to encode the parts into latent codes, which inherit the implicit shape prior to the possible complete object $P$.
@@
 They took the latent codes as input and used graph neural networks to assemble the parts into complete objects.
-They demonstrated their work on 3D chairs, tables, and lamps reassembly tasks.
+They demonstrated their work on 3D chairs, tables, and lamp reassembly tasks.
 Global evaluation as the complete object prior $P$ provides a regularization term to constrain the reassembled results.
@@
 In the image stitching task, an entropy-based metric is proposed to assess the quality of the stitched panoramic image [214].
-In the document reassembly task, Optical Character Recognition (OCR) for word or sentence detection [211, 212].
+In the document reassembly task, Optical Character Recognition (OCR) is used for word or sentence detection [215, 216].
 These methods use pair-wise features to suggest candidate alignments of the fragments.
 They keep the candidate alignments when the evaluation results are promising, while they discard the alignments when the evaluation results are poor.
-Recent approaches used the capacity of large-scale datasets, trained neural networks as discriminators, and used them for global evaluation.
-The discriminator distinguishes the reassembled the results and the natural objects to push the network propose a better reassemble configuration that makes the results look natural.
+More recent approaches leverage large-scale datasets to train neural networks that can evaluate assembly quality.
+These networks learn to distinguish between correct assemblies and incorrect ones, helping guide the reassembly process toward more plausible configurations.
 Li et al.
 [202] propose a GAN to solve the Jigsaw puzzles.
-The designed a network-based discriminator and an adversarial loss to classify the an image is reassembled or real.
-The adversarial loss pushes the GAN to assemble images from pieces that looks natural.
+They designed a network-based discriminator and an adversarial loss to classify the an image is reassembled or real.
+The adversarial loss pushes the GAN to assemble images from pieces that look natural.
 Then a discriminator was trained as adversarial loss to optimize the mating configuration that should look visually realistic like an object.
@@
 (Right) Jigsaw++ [219]: From left to right – input, Jigsaw++ synthetic result, and the ground truth object.
-In summary, local-global alignment often works well on a specific category that has a typical complete object or a large enough database as reference.
-The quality of the initial alignment depends on how to encode the part-global relationship into features.
-Due to the lack of well-annotated database, local-global alignment might be hard to generalize to other object reassembly tasks.
-In contrast, the global evaluation function provides a global shape guidance to regularize the reassembled results.
-However, finding a proper evaluation function for efficient reassembly is still challenging.
+Each approach offers distinct advantages for template-based reassembly.
+Local-global alignment is particularly effective when working with well-defined object categories that have representative complete objects or comprehensive reference databases.
+The success of these methods depends heavily on how effectively they can capture relationships between fragments and complete objects.
+Meanwhile, global evaluation approaches provide flexible guidance for assembly without requiring explicit templates, although the design of effective evaluation metrics remains an active research challenge.
+Future advances may come from combining these complementary approaches.
 6.2 Symmetries
 6.2.1 Symmetry Detection
-Axial symmetry and reflectional symmetry are the two commonly studied symmetry types in geometry analysis and processing, and especially in 3D reassembly.
+Axial symmetry and reflectional symmetry are the two commonly studied symmetry types in geometry analysis and processing, especially in 3D reassembly.
+For a point cloud $P$, symmetry detection aims to find a transformation $s$ (either axial or reflectional) that minimizes:
+$E(s)=\sum_{p\in P}\min_{q\in P}||s(p)-q||^{2}+\lambda R(s)$ (3)
+where $s(p)$ is the transformed point under symmetry transform $s$, $q$ is its closest point in $P$, and $R(s)$ is a regularizer that ensures the transformation is a valid symmetry.
+For axial symmetry, $s$ represents rotation around an axis, while for reflectional symmetry, $s$ represents reflection across a plane.
 As shown in Figure 14 (top), axially symmetric, also referred to as rotationally symmetric, indicates that if the complete object’s exterior surface is intersected by a plane perpendicular to its axis, the resulting cross section is a circle or closely resembles one.
 As a prominent geometrical property, rotational symmetry can serve as a feature to guide the 3D reassembly of fragmented relics, such as sherds.
-Numerous studies explored the estimation of the axis of symmetry.
-Earlier studies estimate the rotational axis using differential geometric properties, such as the sphere of curvature [216] and normal intersection [217].
+Research on symmetry axis estimation has evolved significantly over time.
+Early approaches focused on using differential geometric properties, including the sphere of curvature [220] and normal intersection [221] methods.
 Subsequently, enhancements of the estimated rotational axis have been proposed using M-estimators [222] or Bayesian approaches [223, 224].
-However, these methods can be sensitive to geometric deformation, incompleteness, or noise, often demonstrated in fragmented pieces, leading to inaccurate detection of an object’s symmetrical axis.
+These early methods, while foundational, often struggled with real-world challenges such as geometric deformation, missing pieces, and surface noise - common issues when working with fragmented objects.
 [225] suggested a workflow for axis estimation based on minimum thickness direction.
@@
 [217] showed that the identification of the symmetry axis from surface normal lines may yield two meaningful local minima.
-A multiple-solutions RANSAC algorithm is proposed for initial estimates, followed by a coordinate-descent algorithm to refine these estimates.
+A multiple solution RANSAC algorithm is proposed for initial estimates, followed by a coordinate descent algorithm to refine these estimates.
 The authors showed that their method ensures convergence, achieves a faster convergence rate, and provides state-of-the-art performance in axis estimation.
-As illustrated in Figure 13 (bottom), reflectional symmetry refers to the property that an object does not change upon undergoing a reflection.
-Human skull typically exhibits reflectional symmetry along the mid-sagittal plane [224].
+As illustrated in Figure 14 (bottom), reflection symmetry refers to the property that an object does not change upon undergoing reflection.
+The human skull typically exhibits reflection symmetry along the mid-sagittal plane [228].
 This kind of symmetry is one of the main geometric priors used in skull reassembly and completion [205, 204].
@@
 This method reduces the search space from 6 DOF (for two fragments) to 1 DOF (relative position of two pieces along the axis of symmetry).
+[147] exploits the results of an ongoing reassembly to predict the shape of the resulting object using symmetry-based expansion and then performs alignment of disjoint parts with the predicted shape to also assemble non-contacting parts.
 [227] relaxed the above constraint by encoding the information about the axis as a descriptor into each break line for efficient reassembly.
@@
 Although this approach shows significant potential for reconstructing daily objects, scaling it to handle more diverse or even open-vocabulary input shapes requires larger models and more data.
-The challenge remains to generate shape priors when the underlying shape space exhibits large geometric and topological variations.
-Another limitation is that it does not compute correspondences between the reconstruction and the input fragments.
-Therefore, it does not enforce the rigidity constraints among fragments.
+Significant challenges remain in developing effective shape priors for objects with diverse geometries and topologies.
+Additionally, current methods often lack mechanisms for establishing correspondences between reconstructed shapes and input fragments, making it difficult to enforce rigidity constraints between fragments.
 7 Applications
-We discuss five major application areas of fractured object reassembly (Section 7.1 to Section 7.5) and the relevant application in shape modeling via part assembly.
+We discuss four main applications areas for fractured object reassembly (Section 7.1 to Section 7.3) and the relevant application in shape modeling via part assembly.
 7.1 Bone Reduction
-3D bone reduction serves as a key step in preoperative surgical planning [227, 228, 229] and constitutes a critical element of intraoperative navigation to reduce fractures [230, 231] or a robotic system [232, 233](see Figure 15).
-The objective of bone reduction is to restore the native anatomical configuration of bony structures, encompassing both individual bones and articulations.
-Unlike other fractured entities, fractured bones exhibit the following characteristics.
-Primarily, human bones exhibit pronounced shape priors, which can be leveraged to facilitate the reassembly procedure (see Section 6).
-Owing to the inherent reflection symmetry of human anatomy, the intact contralateral bones can serve as the archetypal reference for the pre-fracture bone structure.
-Secondly, unlike other objects, bone is composed of compact (cortical) bone, cancellous tissue, and cartilage.
-In instances where the fracture occurs within the cancellous tissue, especially within the metaphyseal region of a long bone, the resultant fracture interface commonly presents considerable indeterminacy, which affects the precise alignment of bone fragments [214, 234].
-In the literature, the general workflow of bone reduction includes four components, i.e., CT bone segmentation, template matching, pairwise fragment matching, and multi-fragment matching.
 Figure 16: The surgical robot designed for bone reduction is detailed in [231, 232].
@@
 At the bottom, the figure presents the surgical planning system associated with the robot.
-CT bone segmentation endeavors to delineate the target fractured bone from adjacent healthy bone and to discriminate the bone fragments from each other in CT images.
-Subsequent to delineation, bone fragments are commonly divided into the fracture zone and the intact zone, a process that constitutes one of the most arduous challenges in CT bone segmentation.
-The methods identifying fracture zone are comprised of automatic approaches such as watershed [227], probabilistic watershed [237], template-matching [234] or semi-automatic methods such as region growing [238, 239] or continuous max-flow [230].
-In addition to the non-DL methods mentioned above,  [240] implements supervised contrastive learning with a novel Siamese deep neural network so as to identify and localize the pelvic fracture zones.
-In [234], an approach was introduced to synergistically segment and reassemble the fragments utilizing the contralateral bone as a template and exploiting the synergistic interaction between segmentation and reduction (see Figure 16).
-In recent advancements delineated in [241], an innovative dual-stream learning framework has been introduced for the automated segmentation and categorical labeling of pelvic fractures.
-This framework facilitates the identification of pelvic fracture fragments in varied quantities and anatomical locations through a dual-branch architecture that harnesses distance learning based on bone fragments.
-Considering that the bone template is constituted by the outer surface of the intact zone, aligning the bone fragment with the template necessitates the segmentation of the outer surface of the intact zone [237, 234].
-The bone surface can be extracted by the MarchingCube algorithm [47].
-Template matching is employed as an initial coarse reduction step by aligning the bone fragments with the bone template.
-The healthy contralateral mirrored bone frequently serves as a reference template for the reduction of the fractured side [227, 234, 242, 237, 243, 244].
-Matching the bone fragments with the contralateral bone template can be formulated as a part-to-whole matching problem for 3D rigid objects.
-However, the existence of intra-individual variations, particularly discrepancies in axial torsion, led [245] to conclude that preoperative planning methodologies aimed at reconstructing intricate proximal humerus fractures should not depend uncritically on contralateral anatomical structures.
-Furthermore, contralateral bone is not always available due to ethical considerations in medical practice and instances where the pathology affects both sides.
-Due to the similar morphological characteristics of human bones, the parametric statistical shape model (SSM) serves as an alternative template for fracture reduction [214, 228, 230].
-The alignment of bone fragments with the parametric morphable template is formulated as a part-to-whole problem for 3D non-rigid objects, whereby the output encompasses the rigid 6D pose of each bone fragment and the deformation parameters of the SSM model.
-Moreover, to address the complexities associated with joint dislocation, a statistical pose model (SPM) has been meticulously developed to accurately represent the anatomical pose of the bones comprising the bony joint [246].
-Significantly, the seminal work by [230] integrates both the SSM and the SPM frameworks to address intra-bone fractures and inter-bone dislocations.
-Given the inherently non-convex nature of this complex problem, the iterative procedure alternates through three critical stages: fragment alignment update, SSM adaptation, and SPM adaptation.
+3D bone reduction is essential for preoperative surgical planning [233, 234, 235] and intraoperative navigation [236, 237], including robotic surgery [238, 239](see Figure 16).
+The objective of bone reduction is to restore the anatomical alignment of fractured bones.
+Bone fragments have two unique characteristics that distinguish them from other fractured objects: First, human bones have strong shape priors due to their anatomical nature, including bilateral symmetry which allows the use of contralateral bones as templates (see Section 6).
+Second, bones have complex internal structures including cortical bone, cancellous tissue, and cartilage.
+When fractures occur in cancellous regions, especially in metaphyseal areas of long bones, the irregular fracture surfaces can make precise alignment challenging [218, 240].
+Skull reconstruction, as a specialized case in bone reduction, presents unique characteristics and challenges.
+Unlike long bones, skull fragments are typically modeled as a collection of surfaces, where boundary loops serve as key features for matching [204, 205].
+Feature descriptors computed on these boundary curves enable reliable fragment matching.
+The reconstruction process leverages the skull’s natural bilateral symmetry [228], particularly useful when fragments are missing from one side.
+For cases involving multiple feature curves on a single fragment, the matching problem can be formulated as the Largest Common Point-set (LCP) problem [200].
+Statistical shape models trained on skull datasets provide additional guidance for cases with severe damage [218].
+The precision requirements are particularly stringent due to the skull’s critical role in both surgical planning and forensic analysis.
+The bone reduction workflow typically consists of four steps: CT bone segmentation, template matching, pairwise fragment matching, and multi-fragment matching.
+CT bone segmentation endeavors to separate fractured bones from surrounding healthy tissue and divides individual fragments in CT images.
+A key challenge is distinguishing between fracture zones and intact zones.
+This can be done using automatic methods like watershed [233], probabilistic watershed [241], template-matching [240], or semi-automatic approaches like region growing [242, 243] and continuous max-flow [236].
+Recent deep learning approaches include supervised contrastive learning [244], synergistic segmentation with contralateral bone templates [240] (see Figure 17), and dual-stream learning [245].
+The outer surface of the intact zone [241, 240] serves as a template and can be extracted using the MarchingCube algorithm [48].
+Template matching provides initial coarse alignment between bone fragments and a template.
+Most commonly, the healthy contralateral bone serves as a template [233, 240, 246, 241, 247, 248], though this approach has limitations due to anatomical variations [249] and may not always be available.
+As an alternative, parametric statistical shape models (SSM) [218, 234, 236] can be used.
+For complex cases involving joint dislocations, statistical pose models (SPM) [250] have been developed.
+A notable advancement by [236] combines both SSM and SPM in an iterative procedure that alternates between fragment alignment, SSM adaptation, and SPM adaptation.
 Figure 17: Synergistically segment and reassemble the fragments by utilizing the contralateral bone as a template [240].
-Pair-wise matching aims at aligning two fragments so that their fracture boundaries can perfectly match.
-Pair-wise matching can be divided into two steps: identification of fracture boundary and matching of the fracture boundary.
-After template matching, the fracture boundaries between fragments are close.
-Subsequently, the Euclidean distance between the surface of two fragments can be used as a criterion to identify the fracture boundary [227].
-In instances where fractures occur within rigid regions such as cortical bones, the resulting fracture boundary typically manifests pronounced curvature.
-Subsequent studies [242, 247, 248, 249, 237, 250] have used this characteristic to effectively identify fracture boundaries.
-In contrast, in cases where fractures occur within more flexible regions, such as trabecular bone, the fracture surface tends to appear relatively planar, making the fracture boundary less distinct.
-In such circumstances, interactive annotation of the fracture boundary becomes indispensable [251, 242].
-In the context of matching fracture boundaries, if the fracture occurs within the rigid cortical bone, the methodologies for pair-wise matching bear significant resemblance to those employed in the reassembly of precise artifacts, such as cultural relics.
-These techniques [83, 238, 249] frequently use criteria in which tangential contact points on the respective bone fragments converge at the coordinates of the coincident point, with surface normals oriented oppositely, and matching surface curvatures exhibiting inverse signs.
-Concurrently, the Iterative Closest Point (ICP) algorithm is an effective instrument to execute pairwise matching [252, 242, 248, 250, 227].
-In contrast, in instances where the fracture occurs within the more pliable, trabecular bone, or in scenarios where diminutive fragments are absent near the boundary, pair-wise matching is likely to be ineffective.
-In such circumstances, the previously discussed template matching emerges as an essential and indispensable methodology for bone reduction.
-Note that while common pairwise registrations and pairwise matching techniques such as ICP and feature matching are used in pairwise bone fragment matching, the approaches introduced above instill domain-specific knowledge of individual bones.
-Multi-fragment matching, recast as a quintessential NP-hard problem, seeks to globally optimize the poses of multiple fragments.
-The input for multi-fragment matching includes the current poses of all bone fragments following template matching and pair-wise matching.
-The objective of global optimization is to maximize the region matching score, which combines the scores of template matching (if a template exists) and pair-wise matching [189].
-To achieve this, existing methodologies typically employ an exhaustive search [238] or a greedy search strategy [253] incorporating heuristic rules to address the aforementioned global optimization.
-Within the greedy search strategy, bone fragments from the inactive set are sequentially added to the expanding set of already aligned objects, while ensuring the correct alignment of objects within the active set.
-The next fragment in the inactive set is selected based on the highest score, subject to penetration validity checks.
-In [189], the score was calculated as the weighted sum of the scores from prior template matching and pair-wise matching.
-Furthermore, metrics such as the distance between the fragment and the active set [247], the area of the fracture surface [250], and the size of the bone fragment [227] were also used as priority scoring criteria.
-To avoid entrapment in local minima, multiple fragments can be chosen, rather than a single fragment, to add to the active set using the beam search approach [189].
-A notable challenge in multi-fragment matching is the potential accumulation of pairwise/template errors in multi-piece alignment.
-To address this issue,  [247, 237] introduced the ideal mates concept, which enables the calculation of the multi-piece alignment error without depending on the pairwise registration error [253].
-Unlike [247, 237][189] implemented graph optimization on the target graph, representing the final correspondences among the assembled bone fragments, to globally refine existing transformations.
-Similar to pairwise matching, there are similarities between the multi-fragment matching approaches introduced for bone reduction and those used in restoring stone fragments.
-However, most approaches to bone reduction consider domain-specific knowledge.
-Deep learning in bone reduction.
-In the literature, the deep learning models were trained to segment the cortical surface of bone fragments [234] and fractured bones [254], identify and localize bone fractures [240], and compute bone features [214, 234] for template matching.
-Recently, a significant study [231] introduced an innovative bidirectional framework for automated pelvic fracture surgical planning, which integrates fracture simulation with structural restoration.
-The fracture simulation module specifically considered patient-specific pelvic anatomy, bone density variations, and fracture stochasticity to generate diverse fracture scenarios from intact pelvises.
-For the creation of training datasets that incorporate actual fractures, images of the fractured side and contralateral healthy side were utilized as input and ground-truth data, respectively.
-Accordingly, being trained on both simulated and real fractured bones, the proposed deep structural restoration network reconstructs the bone’s pre-fracture state, subsequently serving as the template for matching the input bone fragments.
-7.2 Fractured Thin-shell Object Reassembly
-In the context of 3D reassembly tasks, a thin-shell object refers to a structure with a thickness that is negligible compared to its other dimensions.
-This implies that the object’s surface geometry, rather than its volume, primarily defines its shape.
-Thin-shell objects present significant challenges during reassembly due to their intricate nature, making the matching and alignment of fragments more complex.
-In forensic skull reassembly, the skull fragments are considered thin shell objects because their thickness is minimal compared to their overall size.
-This requires a precise matching of the fracture regions and intact surfaces to accurately reconstruct the original shape.
-Yin et al.
-[200] and Wei et al.
-[201] model the set of skull fragments as a collection of surfaces.
-They detect the boundary loops of the skull fragments as feature curves and then compute feature descriptors on these curves.
-The feature curves are used for pair-wise matching and alignment computation.
-However, Zhang et al.
-[189] note that due to the thickness of the skull fragments, the feature curves can form multiple loops and may be ambiguous in a single fragment.
-To match multiple feature curves, they formulate it as the Largest Common Point-set (LCP) problem.
-They developed a voting algorithm using 4-point congruent sets extraction to solve this LCP problem.
+Pair-wise matching aligns two fragments by matching their fracture boundaries, involving fracture boundary identification followed by boundary matching.
+Early pioneering work [251] introduced a computational framework combining geometric analysis with medical imaging for pairwise bone fragment registration.
+After template matching brings fragments into rough alignment, fracture boundaries can be identified using Euclidean distance [233].
+For fractures in the cortical bones, pronounced curvature helps to identify boundaries [246, 252, 253, 254, 241, 255].
+However, fractures in trabecular bone often appear planar, requiring manual annotation [256, 246].
+For cortical bone fractures, matching techniques similar to those used in the restoration of cultural artifacts [85, 242, 254] are employed, using criteria like tangential contact points and opposing surface normals.
+ICP is also commonly used [251, 246, 253, 255, 233].
+For trabecular bone fractures or cases with missing fragments, template matching becomes essential.
+Multi-fragment matching optimizes the poses of multiple fragments globally, using poses from template and pair-wise matching as input.
+The goal is to maximize region matching scores [200] through either exhaustive [242] or greedy search [257].
+The greedy search method incorporates heuristic rules for matching and nex-bone selection, where various criteria, including template and pair-wise matching scores [200], fragment-to-set distance [252], fracture surface area [255], and fragment size [233].
+To avoid local minima, beam search can select multiple fragments simultaneously [200].
+Error accumulation is addressed through the ideal mates concept [252, 241, 257] or graph optimization [200].
+Similar to stone fragment matching, bone reduction approaches often incorporate domain-specific knowledge.
+Deep learning models were trained to segment the cortical surface of bone fragments [240] and fractured bones [258], identify and localize bone fractures [244], and compute bone features [218, 240] for template matching.
+A notable advancement [237] introduced a bidirectional framework combining fracture simulation and structural restoration, where the simulation considers patient-specific anatomy and bone properties to generate training data.
+This approach uses both simulated scenarios and real fracture data, where contralateral healthy sides provide ground truth, to train a network that reconstructs pre-fracture states for template matching.
+7.2 Fractured Cultural Heritage Reassembly
+Cultural heritage restoration represents a primary application domain for computational reassembly techniques.
+Reconstructing damaged artifacts preserves historical objects and provides insights into past civilizations and artistic practices.
+We examine four categories of heritage objects: thin-shell artifacts like pottery, stone fragments from sculptures and buildings, planar objects such as maps and inscribed tablets, and larger-scale archaeological site reconstructions.
+Each category presents distinct geometric challenges - from pottery’s axial symmetry to stone’s complex fracture patterns - requiring specialized computational approaches that build upon core reassembly techniques discussed earlier.
 Figure 18: Photo from [90] that shows a project that reconstruct thin-shell objects from its fragments. (Left) fragments. (Right) A digitized fragment.
-Another relevant scenario is archaeological artifact reassembly such as pottery [87, 255], where broken fragments are often thin.
-Cohen et al.
-[87] first digitized the excavated fragments using a Konica 3D scanner (See Figure 17).
-They treated the scanned fragments as surfaces and computed the parabolic contours on them.
-Finally, they used affine-invariant 3D moments to match the contours and align the fragments.
-Their system shows a magnitude improvement in running time compared to manual solutions.
-Angelo et al.
-[256] proposed a recognition pipeline to determine the regions of the 3D scanned fragments into the rim, internal/external wall, and base.
-To reassemble a complete pottery, they designed various descriptors for different regions of the fragments.
-These descriptors help find correspondences and align the fragments.
-Fu et al.
-[257] take curvature and torsion as characteristic descriptors of the curve and search for the candidate matching by the derivative dynamic time warping (DDTW).
-They then design a voting mechanism and ICP method to align the fragments.
-Wang et al.
-[24] observed that due to the thin shell nature of pottery fragments, thin side strip regions often contain many repetitive fractured patterns and lack sufficient unique geometric features.
-They provide a novel contour-based solution to solve batch matching and a boundary-aware ICP method to tackle fine registration.
-They devised an intrinsic shape descriptor to encode a 2D contour to fuse the front and back scans.
-After each fragment’s front and back scans are paired and initially aligned, to address the issue of small overlap and the lack of salient textures and features in the small overlapping region of the two scans, they propose a boundary-based ICP method to ensure robust ICP iteration.
-Their approach demonstrated the ability to scan and reconstruct over 700 fragments in 10 working hours, achieving high reconstruction accuracy.
-7.3 Wall Fragment Reassembly
-Figure 18: The top row shows projects of restoring wall fragments.
+Thin-shell objects.
+The reassembly of thin-shell archaeological artifacts, particularly pottery [90, 259] and ceramics [32, 227], presents unique challenges due to their distinctive geometric properties.
+These objects are characterized by surfaces where thickness is negligible compared to other dimensions, requiring specialized computational approaches.
+Early work by Cohen et al.
+[90] established a foundational pipeline using digitized fragments from 3D scans, computing parabolic contours on surfaces, and employing affine-invariant 3D moments for contour matching and alignment.
+This approach demonstrated significant speed improvements over manual methods.
+For ceramics specifically, researchers have leveraged their inherent axial symmetry properties.
+Karasik and Smilansky [32] developed efficient scanning protocols specifically for ceramic fragments, while Hong et al.
+[227] introduced an incremental 3D reassembly method that exploits axial symmetry for more robust matching.
+Subsequent research enhanced the reconstruction process through various specialized techniques.
+[260] developed a recognition system that segments fragments into distinct regions (rim, walls, and base), designing region-specific descriptors to guide the matching process.
+[261] introduced a curve-based matching approach using curvature and torsion descriptors combined with derivative dynamic time warping (DDTW), incorporating voting mechanisms and ICP refinement.
+A recent breakthrough [24] addressed a critical challenge in thin-shell reassembly - the lack of distinctive geometric features inside regions.
+Their method combines contour-based batch matching with boundary-aware ICP registration, achieving remarkable efficiency by processing over 700 fragments within 10 hours while maintaining high accuracy.
+This progression of research demonstrates the evolution from basic geometric matching to sophisticated systems that can handle large-scale reconstruction tasks.
+Figure 19: The top row shows projects of restoring planar fragments.
 The bottom row shows projects of restoring building fragments.
@@
 (Bottom-right) Reconstruction of the Arch of Titus at the Circus Maximus in Rome [210].
-A special class of thin-shell fragments are wall fragments, in which the complete objects is flat.
-The first well-known effort is Stanford’s Forma Urbis Romae project [258] (Figure 18 (Top-Left)).
-They scanned 1186 fragments that are physically available and, according to estimates, account for about 20% of all fragments.
-As a significant portion of the fragments are missing, they only report dozens of potential matches [79].
-The methods to compute use methods include boundary incision matching, which employ surface textual features, wall feature matching, which is based on location prediction using a underlying map, and multivariate clustering.
-Multivariate clustering finds fragment pairs that may not have explicitly corresponding incisions, but may share other common characteristics including fragment thickness, marble veining direction, axial direction of the incised architecture, presence and orientation of slab edges, and the back surface condition of the fragments (rough, smooth, or unpreserved.
-Another category of notable projects are efforts to reconstruct and match Theran wall paintings [259, 260, 22, 261, 140, 262, 144], including a series of work at Princeton University on reconstruction [22], matching [260, 261, 140], and analysis of fragment boundary patterns for fragment pattern simulation [262].
+Planar heritage.
+Planar heritage objects represent a significant category in computational reassembly, encompassing artifacts such as ancient maps, floor plans, frescoes, and inscribed tablets.
+A landmark project in this domain is Stanford’s Forma Urbis Romae project [262] (Figure 19 (Top-Left)), which aimed to reconstruct an ancient marble map of Rome.
+Despite scanning 1186 fragments (approximately 20% of the estimated total), the project faced significant challenges due to missing pieces.
+Their methodology combined multiple matching approaches: boundary incision matching using surface textural features, location prediction using map features, and multivariate clustering that considered characteristics like marble veining direction and fragment thickness [81].
+The Theran wall paintings project  [263, 264, 22, 265, 144, 266, 148] represents another significant body of work in planar heritage reconstruction, including a series of work at Princeton University on reconstruction [22], matching [264, 265, 144], and analysis of fragment boundary patterns for fragment pattern simulation [266].
 Moreover, [148] presents a generic algorithm for reassembling wall paintings.
@@
 Another consequence of having a large number of fragments is that the global consistency constraint [144, 148] becomes important for multi-fragment matching.
-One relevant scenario is the restoration of ancient papyrus fragments, which share the common characteristics of being flat.
-[263] addresses the reassembly of these fragments using a convolutional neural network in conjunction with a random forest classifier.
-7.4 Fractured Building Reassembly
-Another special class of fragments are fragments of architectural buildings.
-In this case, virtual restoration is necessary as it is very difficult to move the fragments.
-Similar to the setting of wall fragments, the number of fragments is much larger than in the case of bone fragments or thin shell fragments.
-Moreover, the percentages of missing fragments in the building setting, due to the open environment, are even higher than those in the wall painting setting.
-In this case, solving the fractured object reassembly problem strongly rely prior knowledge of the underlying complete object.
-Moreover, user input, i.e., semi-automatic approaches, are necessary to obtain good results.
-One example is the project to reconstruct the Octagon monument in Ephesos [50].
-The project used a semi-automatic approach to virtually restore the original monument from 130 fragments (about 20% of the total number of fragments).
-This restoration is heavily based on prior knowledge of the monument.
-Another relevant project is the reconstruction of the Arch of Titus at the Circus Maximus in Rome [206].
-Again, many fragments are missing, and the virtual reconstruction relies on knowledge about the underlying symmetric structure of the Arch.
-Other notable efforts include the restorations of mixtilinear buildings in Villa Adriana [264], Macedonian Tombs [265], the Maritime Theatre at Hadrian’s Villa [266], and a project on restoring damaged column drums  [267].
+For inscribed tablets, researchers have addressed the challenge of preserving both geometric and textual information.
+[267] developed methods for cuneiform tablet reconstruction considering both wedge impressions and surface geometry, while [268] introduced a system combining geometric and semantic analysis.
+More recent work [269] has used machine learning to analyze cuneiform characters for matching purposes.
+Recent advances in this field include RePAIR dataset [270], which provides both 2D and 3D puzzles from collapsed frescoes, incorporating real-world challenges such as missing pieces and irregular fragment shapes.
+Related work on flat document restoration, such as papyrus reconstruction [271], has leveraged modern machine learning techniques, combining convolutional neural networks with traditional classifiers to address the unique challenges of document reassembly.
+Stone fragments.
+The reassembly of stone fragments encompasses both individual artifacts like sculptures and large-scale architectural remains, presenting unique challenges due to material properties, scale, and frequent missing pieces.
+Early work [53] introduced geometric matching techniques specifically designed for fractured objects, combining patch-based features with global optimization to handle stone fragments effectively.
+For sculptural fragments, [97] developed methods based on original surface features, while [98] introduced robust scoring functions considering gap volume and contact surface properties.
+Recent work [95] proposed specialized methods for stone tool reassembly using contour points and flake surfaces.
+In architectural reconstruction, the challenges scale significantly with fragment size and quantity.
+The Octagon monument project [51], employing techniques from [53] for pairwise matching and multi-piece assembly, demonstrated semi-automatic restoration using only 20% of original fragments.
+The Arch of Titus at the Circus Maximus in Rome reconstruction [210] leveraged architectural symmetry to handle extensive missing pieces.
+Similar approaches have been applied to other significant sites, including the Maritime Theatre [272] and Macedonian Tombs [273].
+For specific architectural elements, such as column drums, [274] developed specialized pattern matching techniques.
+Recent advances [82] address eroded fragments by developing completion methods before matching.
+A key trend across both domains is the integration of computational methods with expert knowledge, particularly crucial when dealing with weathered surfaces or significant missing portions.
+Figure 20: Applications in scientific domains.
+(Top-Left) DNA sequencing [275] (Top-Right) Protein docking [276] (Bottom-Left) Fossil reconstruction of a bird [277].
+(Bottom-Right) Volcano crystal reconstruction [278].
 7.3 Fractured Object Reassembly in Science
@@
 One such example is docking of proteins and DNA Sequencing.
-In [268], in which mitochondrial genes from the unicellular eukaryote Diplonema are systematically fragmented into small pieces that are encoded on separate chromosomes and individually transcribed.
+In [275], which mitochondrial genes from the unicellular eukaryote Diplonema are systematically fragmented into small pieces that are encoded on separate chromosomes and individually transcribed.
 The reassembly problem is to concatenate them into contiguous messenger RNA molecules.
@@
 It is desirable to develop automatic approaches for more accurate and faster reconstructions.
-[272] studied virtual reconstruction of huminin crania to understand human evolution.
+[280] studied the virtual reconstruction of huminin crania to understand human evolution.
 [281] introduced an interactive approach in Blender to restore a dorsal rib and frontal bone of a small-bodied Jurassic-age armored dinosaur from Africa, The digital restoration engaged all modalities of deformation (translation, rotation, scaling, distortion) and reconstruction (fracture infilling, adding missing bone, surface smoothing).
-Figure 19: Applications in scientific domains.
-(Top-Left) DNA sequencing [268] (Top-Right) Protein docking [270] (Bottom-Left) Fossil reconstruction of a bird [271].
-(Bottom-Right) Volcano crystal reconstruction [274].
 One other potential scientific application is the reconstruction of crystals that “blew up” during eruption because of the fast decompression of trapped gases [278].
@@
 Such reconstructions have great implications in geosciences.
+Figure 21: Part Assembly examples. (Top) Object assembly [282]. (Bottom) Mechanical assembly [283].
+Table 3: Comparison of publicly available datasets for fractured object reassembly. Datasets are listed chronologically.
+Datasets Objects/Groups Pieces per Object Total Pieces Breakdown Type Data Type Data Origin Ground Truth TU Wein [53] 7 6-30 101 Real 3D Artifacts Roughly Presious [284] 7 3-30 69 Real 3D Cathedrals Partial BreakingBad [149] 10474 2-100 8442044 Synthetic 3D - FantasticBreaks [285] 150 2 300 Real 3D Household FIRES [24] 15 7-18 123 Real 3D Pottery RePAIR [270] 18 2-44 1070 Real 2D/3D Frescoes Partial
 7.4 Part Assembly
 Part assembly refers to the process of assembling components that are decomposed in a semantically meaningful way, such as furniture components like table legs[201] and mechanical parts like screws[283].
-Although closely related to the fracture assembly problem, part assembly introduces several key differences.
-It generally assumes that each part has an independent semantic function and that each component is complete and functional on its own, unlike the fractured pieces in fracture assembly.
+Although closely related to fracture assembly problem, part assembly introduces several key differences.
+It generally assumes that each part has an independent semantic function and that each component is complete and functional on its own, unlike fractured pieces in fracture assembly.
 Moreover, part assembly allows for duplicated components, and solutions are often non-unique.
@@
 These developments laid the groundwork for later efforts to apply object assembly techniques to 3D asset generation [293, 294, 295].
-Figure 20: Part Assembly examples. (Top) Object assembly [286]. (Bottom) Mechanical assembly [275].
 More recent work in object reassembly has shifted towards estimating the 6-DoF pose of input parts.
@@
 Assemble Them All [283], which included 12,970 assemblies, addressed the challenge of multi-part, multi-scale input data with a disassembly-guided BFS algorithm.
-Figure 21: Datasets that are suitable for deep learning research in fractured object reassembly.
-(Top-Left) The real Fantastic Breaks dataset [295] which combines both isolated fragments (top) and paired complete objects and their fragments (bottom).
-(Top-Right) A real dataset of fragments with ceramic material [24] is captured using the system shown here.
-(Bottom) A synthetic dataset created by performing fracture simulation on a diverse set of objects [145].
 Apart from these two areas, object assembly also emerged in LEGO assembly.
@@
 As stated previously, the major difference is that part assembly focuses on semantic while fracture reassembly focuses on geometry.
-8 Datasets and Software Packages
-We first discuss existing datasets in Section 8.1. We then discuss existing software packages 8.2.
+Table 4: Benchmark performance of published methods on the Breaking Bad Dataset.
+All approaches were initially trained and tested on the everyday object subset.
+They are then fine-tuned and tested on the artifact subset (unless noted with b).
+Testing was conducted on objects containing 2-20 fragments.
+Methods are listed chronologically.
+everyday artifact RMSE(R) MAE(R) RMSE(T) MAE(T) CD PA time RMSE(R) MAE(R) RMSE(T) MAE(T) CD PA degree degree ×$\times 10^{-2}$$\times 10^{-2}$$\times 10^{-3}$ ms degree degree ×$\times 10^{-2}$$\times 10^{-2}$$\times 10^{-3}$ Global [295] 80.7 69.7 14.8 11.8 14.6 21.8 23.7 83.8 71.8 16.6 13.8 19.0 13.3 LSTM [306] 84.2 72.7 16.2 12.7 15.8 19.4 29.9 84.6 73.1 16.8 14.0 21.5 11.7 DGL [212] 79.4 67.8 15.8 12.5 14.3 23.9 28.7 81.7 69.7 16.6 13.8 17.3 19.4 SE(3)-Equiv [146]c 79.3 - 16.9 - 28.5 8.4 129.9 - - - - - - Jigsaw [62] 42.3 36.3 10.7 8.7 13.3 57.3 1063.5 52.4 45.4 22.2 19.3 14.5 45.6 b PHFormer [154] 26.1 22.4 9.3 7.5 9.6 50.7 10.7 32.1 27.7 11.4 9.5 11.8 37.1 DiffAssemble [157] 73.3 - 14.8 - 27.5 - - - - - - - PuzzleFussion++ [145] 38.1 - 8.0 - 6.0 70.6 928.5 52.1 - 13.9 - 14.3 49.6 b PMTR [151] a 31.6 - 9.9 - 5.6 70.6 - 31.6 - 10.1 - 4.3 71.6 FragmentDiff [158] a 13.7 11.5 7.4 5.8 - 90.2 - 18.2 - 8.1 - - 82.3 b
+  • a
+    Reported model is trained and tested on the “volume-constrained” subset.
+  • b
+    Reported model is only trained on everyday subset and is not finetuned on the artifact subset.
+  • c
+    Original paper tested on 2-8 fragments; results for 2-20 fragments are cited from PuzzleFusion++.
+Figure 22: Datasets used for deep learning research in fractured object reassembly.
+(Top-Left) Fantastic Breaks dataset [285] which combines both isolated fragments (top) and paired complete objects and their fragments (bottom).
+(Top-Right) A real dataset of fragments with ceramic material [24] is captured using the system shown here.
+(Bottom-Left) Synthetic dataset created by performing fracture simulation on a diverse set of objects [149].
+(Bottom-Right) RePAIR dataset [270] which contains scans of a fresco at Pompeii archaeological park and ground truth provided by archaeologists.
+8 Datasets and Open Source
+We first discuss existing datasets in Section 8.1 and the evaluation metrics used in benchmarks in Section 8.2. We then discuss existing open-sourced methods in Section 8.3.
 8.1 Datasets
 Machine learning approaches have shown great advancements in the task of reassembly of fragments, but are highly dependent on large-scale datasets.
-However, the first data sets used in fragment reassembly research are primarily small-scale, often manually created [52, 298, 258] or derived from automated semantic segmentation [293, 190].
+However, the first data sets used in fragment reassembly research are primarily small-scale, often manually created [53, 307, 262, 284] or derived from automated semantic segmentation [302, 201].
+Several recent efforts have focused on scanning fragmented objects at scale.
+The first example is Fantastic breaks [285], which includes a small collection of objects with broken fragments and more objects that were broken into fragments by the authors (see Figure 22(Top-Left)).
+Both fragments and complete objects (if they exist) are scanned.
+[24] developed a system to batch scan fragments of objects with ceramic material (see Figure 22(Top-Right)).
+This dataset has ground-truth complete objects for evaluation.
+In addition, the scale of these two datasets makes it possible to develop deep neural networks for reassembly.
+However, the number of categories in these datasets is still small, presenting generalization issues in novel categories.
+The RePAIR dataset [270] (see Figure 22(Bottom-Right)) represents a new advance in real-world fragment assembly datasets, built from a collapsed fresco excavated from archaeological sites.
+It provides both 2D and 3D puzzles along with high-resolution texture information and expert archaeological annotations.
+It includes scenarios with missing pieces, highly irregular fragment shapes, and unknown matching relationships between pieces, which breaks a key assumption present in most existing object-level assembly datasets - that all pieces originate from a single, complete object.
+In addition to real data sets, Breaking Bad [149] is a synthetic data developed by running a simple fracture simulation procedure to generate fragments of shape collections that undergo large geometric and topological variations (see Figure 22(Bottom-Right)).
+This dataset is interesting for fractured object reassembly because the underlying objects belong to diverse object categories.
+For example, template-based reassembling techniques can be developed by learning a shape prior of the complete objects in this dataset.
+Just as the role of ShapeNet [308] in boosting 3D deep learning rich, we envision that research progress on algorithmic reassembly of fractured objects will be accelerated with Breaking Bad and similar datasets.
+However, the fragments are generated using non-physics based fracture simulation.
+The question is how the reassembly pipelines derived from such synthetic datasets generalize to real data.
+We present a comparison of publicly available datasets in Table 3.
 Figure 23: Illustrations of future directions (Left) Representations and algorithms for matching complete template shapes with large geometric and topological variations and end-to-end learning of the entire reassembly pipeline.
 (Middle) Leveraging foundation models to solve sub-tasks such as segmenting fractured surfaces, matching fragments, predicting complete template shapes from input fragments, and matching complete template shapes and fragments.
-(Right) Learning to generate synthetic datasets of 3D fragments and and doing so jointly with learning the reassembly network.
-There are two recent efforts which have focused on scanning fragmented objects at scale.
-The first example is Fantastic breaks [295], which include a small collection of objects with broken fragments and more objects that were broken into fragments by the authors (see Figure 21(Left)).
-Both fragments and complete objects (if they exist) are scanned.
-Another example is [24], which developed a system to batch scan fragments of objects with ceramic material (see Figure 21(Right)).
-This dataset has ground-truth complete objects for evaluation.
-In addition, the scale of these two datasets makes it possible to develop deep neural networks for reassembly.
-However, the number of categories in these two datasets is still small, presenting generalization issues in novel categories.
-In addition to real data sets, Breaking Bad [145] is a synthetic data developed by running a simple fracture simulation procedure to generate fragments of shape collections that undergo large geometric and topological variations ((see Figure 21(Right)).
-This dataset is interesting for fractured object reassembly because the underlying objects belong to diverse object categories.
-For example, template-based reassembling techniques can be developed by learning a shape prior of the complete objects in this dataset.
-Just as the role of ShapeNet [299] in boosting 3D deep learning rich, we envision that research progress on algorithmic fractured object reassembly will be accelerated with Breaking Bad and similar datasets.
-However, as the fragments are generated using non-physics based fracture simulation.
-The question is how the reassembly pipelines derived from such synthetic datasets generalize to real data.
-8.2 Software Packages
+(Right) Learning to generate synthetic datasets of 3D fragments and doing so jointly with learning the reassembly network.
+8.2 Evaluation Metrics
+To quantitatively evaluate reassembly methods, the research community has adopted several standard metrics that measure both the accuracy of predicted transformations and the quality of the final assembly.
+For rotation evaluation, two primary metrics are commonly used: the Mean Absolute Error (MAE) and Root Mean Square Error (RMSE).
+Given a predicted rotation $\tilde{R}$ and ground-truth rotation $R^{gt}$ represented in Euler angles, these metrics are defined as:
+[…]
+Similarly, for translation evaluation, given a predicted translation $\tilde{\boldsymbol{t}}$ and ground-truth translation $\boldsymbol{t}^{gt}$, the corresponding metrics are:
+[…]
+To evaluate the quality of the assembled result, two additional metrics are widely used.
+Chamfer Distance (CD) evaluates the overall assembly quality by measuring the bi-directional point cloud distance between the assembled result and the ground truth shape:
+$\textup{CD}(X,Y)=\frac{1}{|X|}\sum_{x\in X}\min_{y\in Y}||x-y||_{2}+\frac{1}{|Y|}\sum_{y\in Y}\min_{x\in X}||y-x||_{2}$ (6)
+Part Accuracy (PA) measures the percentage of correctly placed fragments. A fragment is considered correctly placed if its Chamfer distance to the ground truth is below a threshold (typically 0.01):
+PA=number of fragments with CD<0.01total number of $\textup{PA}=\frac{\text{number of fragments with CD}<0.01}{\text{total number of fragments}}$ (7)
+Finally, computational efficiency is measured through the total execution time of the algorithm, typically measured in milliseconds.
+This metric is particularly important for practical applications where real-time or near-real-time performance is desired.
+For multi-piece reassembly tasks, the transformation errors (MAE and RMSE) and CD are typically computed as the mean error across all constituent pieces of each object.
+The overall performance of a method is then evaluated by averaging these object-level metrics across the entire test set.
+While these metrics provide a standardized way to compare different reassembly approaches, it is worth noting that they may not fully capture aspects such as physical validity or aesthetic quality of the final assembly.
+8.3 Open Source
 Few early approaches to fractured object reassembly released code.
-A major factor is that in the early days we did not have established benchmark datasets for experimental evaluation.
-For example, there are rich benchmark datasets in computer vision and people deliver open-source codes that execute on benchmark datasets.
+A major factor is that in the early days, we did not have established benchmark datasets for experimental evaluation.
+For example, there are rich benchmark datasets in computer vision and researchers deliver open-source codes that execute on benchmark datasets.
 With the introduction of large-scale benchmark datasets, we see the trend to release codes.
-For example, many papers that evaluate using Breaking Bad have open-source implementations, including Jigsaw [61], SE(3) equivariance [142], DiffAssemble [153], and PuzzleFusion++ [141].
+For example, many papers that evaluate using Breaking Bad have open-source implementations, including Jigsaw [62], SE(3) equivariance [146], PHFormer [154], DiffAssemble [157], PMTR [151], and PuzzleFusion++ [145].
+We provide a comprehensive benchmark performance on these open-sourced methods in Table 4.
+For procedural methods such as ICP on the large-scale dataset, we encourage the readers to refer to the results presented in Neural Shape Mating [203].
 9 Future Directions
@@
 When targeting general object categories that have large topological and geometrical variations, implicit shape representations, which can encode shapes with varying topology, become the only choice.
-However, it is an open problem to properly define correspondences between fragments, which are normally represented as point clouds.
+However, establishing reliable correspondences between implicit representations of complete shapes and the underlying geometric structure of fragments remains an open research challenge.
 Once we have addressed this open problem, it makes sense to study how to integrate all components of the reassembly pipeline in an end-to-end manner, enabling effective training.
@@
 However, this becomes challenging in the case of the fragment reassembly problem.
-First of all, the number of complete objects that have fragments remain sparse.
+First of all, the number of complete objects that have fragments remains sparse.
 In addition, developing software and hardware systems to scan these fragments is very costly.
 There are two interesting topics related to addressing this data issue.
-The first topic is on how to boost the performance of reassembly using vision foundation models.
-Many sub-tasks in the reassembly pipeline, including fracture surface segmentation, establishing correspondences between pairs of fragments, predicting the underlying complete object from fragments, and matching fragments with a complete object, have counterparts as core geometry processing and 3D vision problems.
+The first topic explores boosting reassembly performance using vision foundation models.
+Many subtasks in the reassembly pipeline - including fracture surface segmentation, establishing correspondences between pairs of fragments, predicting the underlying complete object from fragments, and matching fragments with a complete object - have counterparts in core geometry processing and 3D vision problems.
+However, these tasks face unique challenges in the reassembly context.
+For correspondence establishment, current methods primarily rely on geometric matching using surface features or contours, which shows limited generalization ability across different object categories and fracture patterns.
+The assembly problem requires both geometric matching between fragments and semantic matching between fragments and templates - a dual objective that few existing approaches address.
+Additionally, while there has been significant progress in 3D shape generation using implicit representations and neural fields, predicting complete objects from fragments remains challenging.
+Most point cloud generation methods struggle with the diversity of possible completions, especially when handling rotated or partially observed objects.
 Recently, we have seen that 2D foundation models, either through rendered images or lifted 2D features, have led to breakthrough results on these tasks.
-This brings interesting questions on how to adopt these approaches on the corresponding sub-tasks in reassembly.
+2D pretrained model, e.g.
+DINOv2 [309] and Stable Diffusion [310], have shown potential as depth estimation backbone.
+The extracted features have also shown success in finding correspondences [311, 312, 313] with a 30% improvement.
+This brings interesting questions on how to adopt these approaches in the corresponding subtasks in reassembly.
 For example, a recent paper [314] has shown that vision-language models can effectively understand the material properties of objects from rendered images.
-Note that rendering images of fragments is fundamentally different from rendering images of complete objects for visual understanding.
-How to do so to reveal fracture regions is a very interesting research question.
+However, rendering fragments presents unique challenges compared to complete objects, as fracture regions need special consideration for effective visual understanding.
+Developing rendering techniques that can reveal and emphasize fracture characteristics while maintaining other geometric and material properties remains an open research question.
 The second topic is on generating synthetic fracture data to train fracture reassembly networks.
@@
 Such a network allows us to learn hyper-parameters of fracture simulation, e.g., material properties and distribution of external forces, from simulated data to maximize its performance on real data.
-Once we have an end-to-end learnable fracture simulation pipeline, another important problem is to study how to combine fracture reassembly and fracture simulation in an end-to-end manner.
-This allows us to optimize the fracture simulation network, so that the fracture reassembly network generalizes optimally when it is trained from the fracture simulation output.
-This task goes beyond the standard notion of end-to-end training, as we have to develop losses on the output distribution of the fracture simulation network.
+Once we have an end-to-end learnable fracture simulation pipeline, another important problem is studying how to combine fracture reassembly and simulation in an end-to-end manner.
+This allows optimizing the simulation network so the reassembly network generalizes optimally when trained from simulation output.
+This task goes beyond standard end-to-end training, as we must develop losses on the output distribution of the simulation network.
 From a community-building perspective, it is important to create large-scale benchmark datasets for experimental evaluation.
-This will attract more people to work on this problem and, according to the trend in computer vision, will simulate more people to release open-source codes.
+This will attract more researchers to work on this problem and, according to the trend in computer vision, will stimulate more researchers to release open-source codes.
 We do observe a good initial trend from the Breaking Bad dataset [149].
@@
 10 Conclusions
-In this survey, we have provided a thorough survey on the reassembly problem, which finds rich applications in cultural heritage restoration, bone fracture restoration, and object/scene modeling from parts and objects.
-We have covered early non-deep learning based approaches to more recent deep learning based approaches.
-Due to connections to core geometry processing problems in shape segmentation, shape matching, symmetry and primitive detection, and learning shape priors, we also discussed relevant approaches on these problems, emphasizing similarities and differences.
+In this survey, we have provided a thorough survey of the reassembly problem, which finds rich applications in cultural heritage restoration, bone fracture restoration, and object/scene modeling from parts and objects.
+We have covered early procedural approaches to more recent deep learning based approaches.
+Due to connections to core geometry processing problems in shape segmentation, shape matching, symmetry and primitive detection, and learning shape priors, we also discussed relevant approaches to these problems, emphasizing similarities and differences.
 We also covered current state-of-the-art methods for solving real problems in these application domains.
@@
 We conclude that this very fundamental research field can also benefit from having more scientific applications.
-Acknowledgement.
+Acknowledgement
 Qixing Huang is supported by NSF Career IIS-2047677, NSF IIS-2413161, and Gifts from Adobe and Google.
-Xin Li is partly supported by NSF CBET-2115405 and NIH R15HD108765.
+Xin Li is partly supported by NSF-IIS 1320959 and NSF CBET-2115405.
 Junfeng Jiang would like to acknowledge support from Jiangsu Province Key Research Program-Social Development-Clinical Frontier Technologies (BE2022718).